### v3

# CARLA: Self-supervised Contrastive Representation Learning for Time Series Anomaly Detection

Abstract
One main challenge in time series anomaly detection (TSAD) is the lack of labelled data in many real-life scenarios. Most of the existing anomaly detection methods focus on learning the normal behaviour of unlabelled time series in an unsupervised manner. The normal boundary is often defined tightly, resulting in slight deviations being classified as anomalies, consequently leading to a high false positive rate and a limited ability to generalise normal patterns. To address this, we introduce a novel end-to-end self-supervised ContrAstive Representation Learning approach for time series Anomaly detection (CARLA). While existing contrastive learning methods assume that augmented time series windows are positive samples and temporally distant windows are negative samples, we argue that these assumptions are limited as augmentation of time series can transform them to negative samples, and a temporally distant window can represent a positive sample. Existing approaches to contrastive learning for time series have directly copied methods developed for image analysis. We argue that these methods do not transfer well. Instead, our contrastive approach leverages existing generic knowledge about time series anomalies and injects various types of anomalies as negative samples. Therefore, CARLA not only learns normal behaviour but also learns deviations indicating anomalies. It creates similar representations for temporally close windows and distinct ones for anomalies. Additionally, it leverages the information about representations’ neighbours through a self-supervised approach to classify windows based on their nearest/furthest neighbours to further enhance the performance of anomaly detection. In extensive tests on seven major real-world time series anomaly detection datasets, CARLA shows superior performance (F1 and AU-PR) over state-of-the-art self-supervised, semi-supervised, and unsupervised TSAD methods for univariate time series and multivariate time series. Our research highlights the immense potential of contrastive representation learning in advancing the time series anomaly detection field, thus paving the way for novel applications and in-depth exploration in the domain.
keywords:
Anomaly Detection, Time Series, Deep Learning, Contrastive Learning, Representation Learning, Self-supervised learning
[…]
[inst1]organization=Monash University,city=Melbourne, state=Victoria, country=Australia
[…]
[inst2]organization=Griffith University, city=Gold Coast, state=Queensland, country=Australia
[…]
[inst3]organization=IBM T. J. Watson Research Center, city=Yorktown Heights, state=NY, country=USA
1 Introduction
In many modern applications, data analysis is required to identify and remove anomalies (a.k.a. outliers) to ensure system reliability. Several machine learning algorithms are well-suited for detecting these outliers Domingues et al. (2018). In time series data, anomalies can result from different factors, including equipment failure, sensor malfunction, human error, and human intervention. Detecting anomalies in time series data has numerous real-world uses, including monitoring equipment for malfunctions, detecting unusual patterns in IoT sensor data, enhancing the reliability of computer programs and cloud systems, observing patients’ health metrics, and pinpointing cyber threats. Time series anomaly detection (TSAD) has been the subject of decades of intensive research, with numerous approaches proposed to address the challenge of spotting rare and unexpected events in complex and noisy data. Statistical methods have been developed to monitor and identify abnormal behaviour Domingues et al. (2018). Recent advancements in deep learning techniques have effectively tackled various anomaly detection problems Pang et al. (2021). Specifically, for time series with complex nonlinear temporal dynamics, deep learning methods have demonstrated remarkable performance Schmidl et al. (2022).
Most of these models focus on learning the normal behaviour of data from an unlabelled dataset and, therefore, can potentially predict samples that deviate from the normal behaviour as anomalies. The lack of labelled data in real-world scenarios makes it difficult for models to learn the difference between normal and anomalous behaviours. The normal boundary is often tightly defined, which can result in slight deviations being classified as anomalies. Models that cannot discriminate between normal and anomaly classes may predict normal samples as anomalous, leading to high false positives. Figure 0(a) shows an example of a histogram of anomaly scores of an existing unsupervised TSAD method called THOC Shen et al. (2020) on a benchmark dataset (M-6 test set from the MSL dataset). Frequent false positives often occur when normal data are assigned high anomaly scores. In instances where the representation of a segment of the normal data closely resembles that of anomalous data, there can be an elevated anomaly score attributed to these normal samples, consequently leading to an increase in false positives. This issue can be seen in right side of Figure 0(a) where many normal samples are assigned high anomaly scores.
An alternative approach is to leverage self-supervised contrastive representation learning. Contrastive learning entails training a model to differentiate between pairs of similar and dissimilar samples in a dataset. Contrastive representation learning has been successful in image Chen et al. (2020) and natural language processing Logeswaran and Lee (2018), and its potential in TSAD has been explored in recent years. Contrastive loss functions can increase the distance between normal and their corresponding anomalous samples while decreasing the distance between normal samples. By doing so, the distinction between normal and anomalous data is made clearer and more pronounced, which helps to identify a more accurate boundary for the normal samples. However, existing contrastive learning methods for TSAD mainly assume that augmented time series windows are positive samples and temporally distant windows are negative samples. We argue that these assumptions carry the risk that augmentation of time series can transform them to negative samples, and a temporally distant window can represent a positive sample, leading to ineffective anomaly detection performance. Figure 0(b) shows anomaly scores of a contrastive learning TSAD method called TS2Vec Yue et al. (2022). As demonstrated, the anomaly scores of both normal and anomalous data are blended together, resulting in the inefficacy of anomaly detection. In fact, in situations where normal and anomalous data are intermingled within the representation space, achieving an enhanced anomaly detection rate—as measured by metrics such as the F1 score or the area under the precision-recall curve (AU-PR)—is accompanied by an increase in the false positive rate.
Figure 1: Histogram of anomaly scores produced by (a) THOC model Shen et al. (2020) and (b) TS2Vec model Yue et al. (2022). Histograms illustrating the distribution of anomaly scores for the THOC and TS2Vec models using data from entity M-6 of the MSL dataset Hundman et al. (2018).
We propose a novel two-stage framework called CARLA, designed specifically to enhance time series anomaly detection. Our novel approach addresses the lack of labelled data through a contrastive approach which leverages existing generic knowledge about different types of time series anomalies in the first stage (pretext). We inject various types of anomalies, which facilitates the learning representations of normal behaviour. This is achieved by encouraging the learning of similar representations for windows that are temporally closed windows while ensuring dissimilar representations for windows and their corresponding injected anomalous windows.
Additionally, to ensure that the representation of existing real anomalous windows (for which we do not have labels) is different from normal representations, we employ a self-supervised approach to classify normal/anomalous representations of windows based on their nearest/furthest neighbours in the representation space in the second stage (self-supervised classification). By making the normal representations more discriminative, we enhance the anomaly detection performance in our proposed model. Specifically, the main contributions of this paper can be summarised as follows:
  • 1.
    We propose a novel contrastive representation learning model to detect anomalies in time series, which delivers top-tier outcomes across a range of real-world benchmark datasets, encompassing both univariate time series (UTS) and multivariate time series (MTS). Addressing the challenge of lack of labelled data, our model learns to effectively discriminative normal patterns from anomalous in the feature representation space.
  • 2.
    We propose an effective contrastive method for TSAD to learn feature representations for a pretext task by leveraging existing generic knowledge about time series anomalies (see Figure 5(a)).
  • 3.
    We propose a self-supervised classification method that leverages the representations learned in the pretext stage to classify time series windows. Our goal is to classify each sample by utilising its neighbours in the representation space, which were learned during the pretext stage (see Figure 5(b)).
  • 4.
    Our comprehensive analysis across seven real-world benchmark datasets reveals the superior performance of CARLA over a range of ten SOTA unsupervised, semi-supervised, and self-supervised contrastive learning models. CARLA’s consistent balance between False Positive Rate (FPR) and Area Under the Precision-Recall Curve (AU-PR) throughout various MTS and UTS datasets underscores its precision and reliability. This balance ensures that CARLA provides reliable and precise alerts, which are crucial for many real-world applications.
2 Related Work
In this section, we provide a brief survey of the existing literature that intersects with our research. We concentrate on three areas: deep learning for anomaly detection in time series, unsupervised representation of time series, and the contrastive representation learning technique.
2.1 Time Series Anomaly Detection
The detection of anomalies within time series data has been the subject of extensive research, using an array of techniques from statistical methods to classical machine learning and, more recently, deep learning models Schmidl et al. (2022). Established statistical techniques such as moving averages like the ARIMA model Box and Pierce (1970) have seen widespread application. Machine learning techniques, including clustering algorithms and density-based approaches, alongside algorithms similar to decision trees Fei Tony Liu and Zhou (2008) have also been leveraged.
Deep Learning methods in TSAD: In recent years, deep learning has proven highly effective due to its ability to autonomously extract features Pang et al. (2021). The focus within TSAD is largely on unsupervised Hundman et al. (2018), semi-supervised Park et al. (2018), and recently self-supervised Yue et al. (2022) approaches, addressing the issue of scarce labelled data. Techniques such as OmniAnomaly Su et al. (2019) prove to be particularly useful in situations where anomaly labels are unavailable, while semi-supervised methods make efficient use of the labels that are available.
Deep learning techniques, encompassing autoencoders Sakurada and Yairi (2014); Yao et al. (2023), VAEs Park et al. (2018), RNNs Su et al. (2019), LSTM networks Hundman et al. (2018), GANs Li et al. (2019), and Transformers Tuli et al. (2022); Xu et al. (2021a) have shown potential in spotting anomalies within time series, particularly for high-dimensional or non-linear data. The preference for deep models like LSTM-VAE Park et al. (2018), DITAN Giannoulis et al. (2023) and THOC Shen et al. (2020) is because they excel at minimising forecasting errors while capturing time series data’s temporal dependencies.
In summary, semi-supervised methods, such as LSTM-VAE Park et al. (2018) excel when labels are readily available. On the other hand, unsupervised methods, such as OmniAnomaly Su et al. (2019) and AnomalyTransformer Xu et al. (2021a), become more suitable when obtaining anomaly labels is a challenge. These unsupervised deep learning methods are preferred due to their capability to learn robust representations without requiring labelled data. The advent of self-supervised learning-based methods has further improved generalization ability in unsupervised anomaly detection Zhao et al. (2020).
2.2 Unsupervised Time Series Representation
The demonstration of substantial performance by unsupervised representation learning across a wide range of fields, such as computer vision Chen et al. (2020); Milbich et al. (2020), natural language processing Logeswaran and Lee (2018), and speech recognition Xu et al. (2021b), has made it a desirable technique. In the realm of time series, methods like TKAE Bianchi et al. (2019) and TST Zerveas et al. (2021) have been suggested. While these methods have made significant contributions, some of them face challenges with scalability for exceptionally long time series or encounter difficulties in modelling complex time series. To overcome these limitations, methods such as TNC Tonekaboni et al. (2021) and T-Loss Franceschi et al. (2019) have been proposed. They leverage time-based negative sampling, triplet loss, and local smoothness of signals for the purpose of learning multivariate time series representations that are scalable. Despite their merits, these methods often limit their universality by learning representations of particular semantic tiers, relying on significant assumptions regarding invariance during transformations. Recent studies, such as RoSAS Xu et al. (2023), have advanced outlier detection in datasets beyond time series by utilising triplet loss.
2.3 Contrastive Representation Learning
Contrastive representation learning creates an embedding space where similar samples are close and dissimilar ones are distant, applicable in domains like natural language processing, computer vision, and time series anomaly detection Le-Khac et al. (2020). Traditional models like InfoNCE loss Oord et al. (2018), SimCLR Chen et al. (2020), and MoCo He et al. (2020) use positive-negative pairs to optimize this space, demonstrating strong performance and paving the way for advanced techniques.
In time series anomaly detection, contrastive learning is crucial for recognising patterns. TS2Vec Yue et al. (2022) uses this approach hierarchically for multi-level semantic representation, while DCdetector Yang et al. (2023) introduces a dual attention asymmetric design with pure contrastive loss, enabling permutation invariant representations.
3 CARLA
Problem definition: Given a time series $\mathcal{D}$ which is partitioned into $m$ overlapping time series windows $\{w_{1},..,w_{i},..,w_{m}\}$ with stride 1 where $w_{i}=\{x_{1},..,x_{i},..,x_{WS}\}$, $WS$ is time series window size, $x_{i}\in\mathbb{R}^{Dim}$ and $Dim$ is the dimension of time series, the goal is to detect anomalies in time series windows.
CARLA (A Self-supervised ContrAstive Representation Learning Approach for time series Anomaly detection) is built on several key components, each of which plays a critical role in achieving effective representation learning as illustrated in Figure 2. CARLA consists of two main stages: the Pretext Stage and the Self-supervised Classification Stage.
Initially, in the Pretext Stage (section 3.2), it employs anomaly injection to learn similar representations for temporally proximate windows and distinct representations for windows and their equivalent anomalous windows (section 3.1). These injected anomalies include point anomalies, such as sudden spikes, and subsequence anomalies, like unexpected pattern shifts. This technique not only aids in training the model to recognise deviations from the norm but also strengthens its ability to generalise across various types of anomalies. At the end of the Pretext Stage, we establish a prior by finding the nearest and furthest neighbours for each window representation for the next stage. The Self-supervised Classification Stage (section 3.3) then classifies these window representations as normal or anomalous based on the proximity of their neighbours in the representation space (section 3.4). This classification aims to group similar time series windows together while distinctly separating them from dissimilar ones. The effectiveness of this stage is pivotal in accurately categorising time series windows, reinforcing CARLA’s capability to differentiate between normal and anomalous patterns. The comprehensive end-to-end pipeline of CARLA, including these stages, is illustrated in Figure 2.
The following sections present the ideas underlying our approach.
Figure 2: The end-to-end pipeline of CARLA consists of two main stages: the Pretext Stage and the Self-supervised Classification Stage. In the Pretext Stage, anomaly injection techniques are used for self-supervised learning. The Self-supervised Classification Stage integrates the learned representations for a contrastive approach that maximises the similarity between anchor samples and their nearest neighbours while minimising the similarity between anchor samples and their furthest neighbours. The output is a trained model and the majority class, enabling inference for anomaly detection.
3.1 Anomaly Injection
The technique of anomaly injection, a powerful data augmentation strategy for time series, facilitates the application of self-supervised learning in the absence of ground-truth labels. This technique has recently been applied in TSAD, notably in COUTA Xu et al. (2022). While the augmentation methods we employ are not designed to represent every conceivable anomaly type Xu et al. (2021a) — a goal that would be unattainable — they amalgamate various robust and generic heuristics to effectively identify prevalent out-of-distribution instances.
3.1.1 Anomaly Injection Steps
During the training phase, each window is manipulated by randomly choosing instances within a given window $w_{i}$. Two primary categories of anomaly injection models—point anomalies and subsequence anomalies—are adopted to inject anomalies to a time series window $w_{i}$. In a multivariate time series context, a random start time and a subset of dimension(s) $d$ are selected for the injection of a point or subsequence anomalies. Note in the context of multivariate time series, anomalies are not always present across all dimensions, prompting us to randomly select a subset of dimensions for the induced anomalies ($d<\lceil Dim/10\rceil$). The injected anomaly portion for each dimension varied from 1 data point to 90% of the window length.
This approach fosters the creation of a more diverse set of anomalies, enhancing our model’s capability of detecting anomalies that exist in multiple dimensions. This approach strengthens our model’s effectiveness in discriminating between normal and anomalous representations in the Pretext Stage. It is important to underscore that the anomaly injection strategies employed in our model’s evaluation were consistently applied across all benchmarks to ensure a fair and impartial comparison of the model’s performance across various datasets. The detailed step-by-step process for our anomaly injection approach is elaborated in Algorithm 1. This comprehensive algorithm encapsulates the methodology of both point anomaly injection and subsequence anomaly injection, providing a clear understanding of our process and its implementation.
Algorithm 1 InjectAnomaly($w$)
1:Time series window $w$
2:Anomaly injected time series window $w^{\prime}$
3:$types\leftarrow\{\texttt{Seasonal},\texttt{Trend},\texttt{Global},\texttt{Contextual},\texttt{Shapelet}\}$
4:$Dim\leftarrow\text{random subset of dimensions of }w~{}\text{(for UTS, }Dim=\{1\}\text{)}$
5:$w^{\prime}\leftarrow\text{copy of}~{}w$
6:$s,e\leftarrow\text{random start and end points from }(0,size(w)]\text{, where }e>s$
7:for each dimension $d$ in $Dim$ do
8:    $anomaly\leftarrow\text{randomly choose anomaly type from }types$
9:    if $anomaly=\texttt{Global}$ then
10:        $\mu_{w},\sigma_{w}\leftarrow\text{mean and std of }w\text{ in dimension }d$
11:        $g\leftarrow\text{random number form }[3,5]$triangleright coefficient for Global anomaly
12:        $w^{\prime}(s)=\text{random value from }\{\mu_{w}+g.\sigma_{w},\mu_{w}-g.\sigma_{w}\}$
13:    else if $anomaly=\texttt{Contextual}$ then
14:        $\mu_{se},\sigma_{se}\leftarrow\text{mean and std of subsequence }w(t|t\in[s,e])\text{ in dimension }d$
15:        $x\leftarrow\text{random number form }[3,5]$triangleright coefficient for Contextual anomaly
16:        $w^{\prime}(s)=\text{random value from }\{\mu_{se}+x.\sigma_{se},\mu_{se}-x.\sigma_{se}\}$
17:    else if $anomaly=\texttt{Seasonal}$ then
18:        $f\leftarrow\text{random number from }\{\frac{1}{3},\frac{1}{2},2,3\}$triangleright frequency coefficient for Seasonal anomaly
19:        $w^{\prime}(t)=\begin{cases}w(s+\left(\left\lfloor(t-s)\cdot k\right\rfloor\mod n\right))&\text{if }s\leq t<e\text{ and }k>1\\ w(s+\left\lfloor(t-s)\cdot k\right\rfloor)&\text{if }s\leq t<e\text{ and }0<k<1\\ w(t)&\text{otherwise}\end{cases}$
20:    else if $anomaly=\texttt{Trend}$ then
21:        $b\leftarrow\text{random number form }[3,5]$triangleright coefficient for Trend anomaly
22:        $w^{\prime}(t)=\begin{cases}w(t)+b.\sigma_{w}&\text{if }s\leq t\leq e\\ w(t)&\text{if }t<s\text{ or }t>e\end{cases}$
23:    else if $anomaly=\texttt{Shapelet}$ then
24:        $w^{\prime}(t)=\begin{cases}w(s)&\text{if }s\leq t\leq e\\ w(t)&\text{if }t<s\text{ or }t>e\end{cases}$
25:    end if
26:end for
27:Return $w^{\prime}$
3.1.2 Point Anomalies
Our methodology incorporates the injection of single-point anomalies (a spike) at randomly chosen instances within a given window $w_{i}$. We employ two distinct types of point anomalies, namely, Global (Figure 2(b)) and Contextual (Figure 2(c)Xu et al. (2021a).
3.1.3 Subsequence Anomalies
The technique of subsequence anomaly injection, motivated by the successful implementation of the Outlier Exposure approach Hendrycks et al. (2019), enhances anomaly detection in time series data by generating contextual out-of-distribution examples. It involves the introduction of a subsequence anomaly within a time series window $w_{i}$, represented as $w(t|t\in[s,e])$, where $s$ and $e$ denotes the anomaly’s starting and end time points. We explore three primary types of subsequence (a.k.a. pattern) anomalies: Seasonal (Figure 2(d)), Shapelet (Figure 2(e)), and Trend (Figure 2(f)Xu et al. (2021a).
(a) Original Window
(b) Global Anomaly
(c) Contextual Anomaly
(d) Seasonal Anomaly
(e) Shapelet Anomaly
(f) Trend Anomaly
Figure 3: Different types of synthetic anomaly injection used in CARLA. The figure presents the effect of synthetic anomaly injections into a randomly selected window of size 400 from the first dimension of entity E-2 in the MSL dataset (Hundman et al., 2018). The (a) represents the original time series window, while the remaining five demonstrate the same window but with different types of anomalies injected. The anomalies are categorised into (b) Global, (c) Contextual, (d) Seasonal, (e) Shapelet, and (f) Trend. In each of the five plots with anomalies, the anomalous points or subsequences are accentuated with a red colour. The synthetic anomalies were introduced using one or more randomly selected anomaly injection methods, applied sequentially across one or more dimensions of the multivariate time series.
3.2 Pretext Stage
The Pretext Stage of CARLA consists of two parts: Part One is a contrastive representation learning step that uses a ResNet architecture (which has shown effectiveness in times series classification tasks Ismail Fawaz et al. (2019)) to learn representations for time series windows, and Part Two is a post-processing step that uses the learned representations from Part One to identify semantically meaningful nearest and furthest neighbours for each learned window representation. Algorithm 2 shows the Pretext steps.
Algorithm 2 PretextCARLA($\mathcal{D}$, $Q$, $\alpha$)
1:Sequential time series windows $\mathcal{D}=\{w_{1},w_{2},...,w_{m}\}$, number of nearest/furthest neighbours $Q$, margin $\alpha$.
2:Trained model $\phi_{p}$, all neighbours set $\mathcal{B}$, nearest neighbours set $\mathcal{N}$, furthest neighbours set $\mathcal{F}$.
3:$\mathcal{T}\leftarrow\varnothing$, $\mathcal{B}\leftarrow\varnothing$, $\mathcal{N}\leftarrow\varnothing$, $\mathcal{F}\leftarrow\varnothing$
4:for $i\leftarrow 1$ to $|\mathcal{D}|$ do
5:    $a_{i}\leftarrow w_{i}$triangleright anchor
6:    $p_{i}\leftarrow w_{i-r}$, where $r\sim\mathcal{U}(1,y)$triangleright positive pair
7:    $n_{i}\leftarrow$ InjectAnomaly($w_{i}$) $\triangleright$ negative pair (Alg. 1)
8:    append triplet $(a_{i},p_{i},n_{i})$ to $\mathcal{T}$triangleright triplets batches
9:    add $a_{i}$ and $n_{i}$ to $\mathcal{B}$triangleright neighbours set with size $2|\mathcal{D}|$
10:end for
11:while Pretext loss $\mathcal{L}_{pretext}(\phi_{p},\mathcal{T},\alpha)$ decreases do
12:    Update $\phi_{p}$ with $\mathcal{L}_{pretext}$triangleright i.e. Equation (1)
13:end while
14:for $j\leftarrow 1$ to $|\mathcal{B}|$ do
15:    $\mathcal{N}_{j}\leftarrow Q$ nearest neighbours of $w_{j}\in\mathcal{B}$ in $\phi_{p}(\mathcal{B})$ space
16:    $\mathcal{F}_{j}\leftarrow Q$ furthest neighbours of $w_{j}\in\mathcal{B}$ in $\phi_{p}(\mathcal{B})$ space
17:end for
18:Return $\phi_{p}$, $\mathcal{B}$, $\mathcal{N}$, $\mathcal{F}$triangleright inputs of the next stage
3.2.1 Part One: Contrastive Representation Learning
In this part, we introduce a contrastive learning framework for learning a discriminative representation of features for time series windows. To extract features from the time series data, we utilise a multi-channel ResNet architecture, where each channel represents a different time series dimension. Using different kernel sizes in ResNet allows us to capture features at various temporal scales, which is particularly important in analysing time series data and makes our model less sensitive to window size selection. We add an MLP layer as the final layer to produce a feature vector with lower dimensions.
To encourage the model to distinguish between different time series windows, we utilise a triplet loss function. Specifically, we create triplets of samples in the form of $(a,p,n)$, where $a$ is the anchor sample (i.e. the original time series window $w_{i}$), $p$ is a positive sample (i.e. another random window $w_{i-r}$, selected from $y$ previous windows, where $r\sim\mathcal{U}(1,y)$), and $n$ is a negative sample (i.e. an anomaly injected version of the original window $w_{i}$ to make it anomalous). Assuming we have all the triplets of $(a,p,n)$ in a set $\mathcal{T}$, the pretext triplet loss function is defined as follows:
$\mathcal{L}_{Pretext}(\phi_{p},\mathcal{T},\alpha)=\frac{1}{|\mathcal{T}|}\sum_{(a,p,n)\in\mathcal{T}}\max(\lVert\phi_{p}(a)-\phi_{p}(p)\rVert_{2}^{2}-\lVert\phi_{p}(a)-\phi_{p}(n)\rVert_{2}^{2}+\alpha,0)$ (1)
Where $\phi_{p}(\cdot)$ is the learned feature representation neural network, $\lVert\cdot\rVert_{2}^{2}$ denotes the squared Euclidean distance, and $\alpha$ is a margin that controls the minimum distance between positive and negative samples. The objective of the pretext triplet loss function is to decrease the distance between the anchor sample and its corresponding positive sample while simultaneously increasing the distance between the anchor sample and negative samples. This approach encourages the model to learn a representation that can differentiate between normal and anomalous time series windows.
Our approach empowers the model to learn similar feature representations for temporally proximate windows. Since the majority of data is normal, the model captures temporal relationships of normal data through learning similar representations for normal windows that are temporally proximate. Furthermore, by introducing anomalies into the system, the model can learn a more effective decision boundary, resulting in a reduced false positive rate and enhanced precision in identifying anomalies when compared to current state-of-the-art models.
The output of Part One includes a trained neural network (ResNet) $\phi_{p}$ and a list of anchor and negative samples stored in $\mathcal{B}$.
3.2.2 Part Two: Nearest and Furthest Neighbours
Part Two of our approach uses the feature representations learned in Part One to identify semantically meaningful nearest and furthest neighbours for each sample. To achieve this, we utilise $\mathcal{B}$ along with the indices of $Q$ nearest and $Q$ furthest neighbours of all samples in $\mathcal{B}$.
The primary goal of this part is to generate a prior that captures the semantic similarity and semantic dissimilarity between windows’ representations using their neighbours, as our empirical analysis shows that in the majority of cases, these nearest neighbours belong to the same class (see Figure 5(a)). Semantic similarity for a given window $w_{i}$ defines as $Q$ nearest neighbours of $\phi_{p}(w_{i})\in\phi_{p}(\mathcal{B})$, where $Q$ is the number of nearest neighbours. And, semantic dissimilarity for a given window $w_{i}$ defined as $Q$ furthest neighbours of $\phi_{p}(w_{i})\in\phi_{p}(\mathcal{B})$. The output of Part Two is a set of all anchor and negative samples (anomaly injected) and the indices of their $Q$ nearest neighbours $\mathcal{N}$ and $Q$ furthest neighbours $\mathcal{F}$. Utilising $\mathcal{N}$ and $\mathcal{F}$ can enhance the performance of our classification method used in the Self-supervised Classification Stage. The culmination of the Pretext stage is a comprehensive set of time series windows alongside their nearest and furthest neighbours, setting the stage for their utilisation in the forthcoming Self-supervised Classification Stage.
3.3 Self-supervised Classification Stage
As we transition into the Self-supervised Classification Stage, we utilise the output of the Pretext stage as our foundational input. This stage includes initialising a new ResNet architecture with the learned feature representations from Part One of the Pretext Stage and then integrating the semantically meaningful nearest and furthest neighbours from Part Two as a prior ($\mathcal{N}$ and $\mathcal{F}$) into a learnable approach. Algorithm 3 shows the steps of the Self-supervised classification Stage.
Algorithm 3 SelfSupevisedCARLA($\phi_{p}$, $\mathcal{B}$, $\mathcal{N}$, $\mathcal{F}$, $C$, $\beta$)
1:Initial trained neural network (ResNet) from the Pretext Stage $\phi_{p}$, Dataset of time series windows including original windows and anomaly injected windows $\mathcal{B}$, Set of $Q$ nearest neighbours for each window $\mathcal{N}$, Set of $Q$ furthest neighbours for each window $\mathcal{F}$, Number of Classes $C$, Entropy loss weight $\beta$
2:Trained model $\phi_{s}$, Majority class $C_{m}$
3:$\phi_{s}\leftarrow$ initialise by $\phi_{p}$triangleright $\phi_{p}$ from Algorithm 2 (PretextCARLA)
4:while $\mathcal{L}_{Self-supervised}(\phi_{s},\mathcal{B},\mathcal{N},\mathcal{F},\mathcal{C},\beta)$ decreases do
5:    Update $\phi_{s}$ with $\mathcal{L}_{Self-supervised}$triangleright i.e. Equation (6)
6:end while
7:for $i\leftarrow 1$ to $|\mathcal{D}|$ do
8:    $C^{i}=\arg\max(\phi_{s}(w_{i})),w_{i}\in\mathcal{D}$triangleright assign class label $C_{j}\in\mathcal{C}$ to window $i$
9:end for
10:$C_{m}=\arg\max_{C_{j}\in\mathcal{C}}(n(C_{j}))$triangleright find majority class $C_{m}$triangleright $n(C_{j})$ denotes the number of members in a class $C_{j}$
11:Return $\phi_{s}$, $C_{m}$
To encourage the model to produce both consistent and discriminative predictions, we employ a contrastive approach with a customised loss function. Specifically, the loss function maximises the similarity between each window representation and its nearest neighbours while minimising the similarity between each window representation and its furthest neighbours. The loss function can be defined as follows: At the beginning of this stage, we have $\mathcal{B}$ from the Pretext Stage, which is the set of all original window representations (we call them anchors in this stage) and their corresponding anomalous representations. Let $C$ be the number of classes. We also have the $Q$ nearest neighbours of the anchor samples $\mathcal{N}$ and the $Q$ furthest neighbours of the anchor samples $\mathcal{F}$.
We aim to learn a classification neural network function $\phi_{s}$ initialised by $\phi_{p}$ from the Pretext Stage - that classifies a window $w_{i}$ and its $Q$ nearest neighbours to the same class, and $w_{i}$ and its $Q$ furthest neighbours to different classes. The neural network $\phi_{s}$ terminates in a softmax function to perform a soft assignment over the classes $\mathcal{C}=\{1,...,C\}$, with $\phi_{s}(w)\in[0,1]^{C}$.
To encourage similarity between the anchor samples and their nearest neighbours, we compute the pairwise similarity between the probability distributions of the anchor and its neighbours as it is shown in Equation (2). The dot product between an anchor and its neighbour will be maximal when the output of the softmax for them is close to 1 or 0 and consistent in that it is assigned to the same class. Then, we define a consistency loss in Equation (3) using the binary cross entropy to maximise the similarity between the anchor and nearest neighbours.
$\text{similarity}(\phi_{s},w_{i},w_{j})=\langle\phi_{s}(w_{i})\cdot\phi_{s}(w_{j})\rangle=\phi_{s}(w_{i})^{\top}\phi_{s}(w_{j})$ (2)
$\mathcal{L}_{consistency}(\phi_{s},\mathcal{B},\mathcal{N})=-\frac{1}{\lvert\mathcal{B}\rvert}\sum_{w\in\mathcal{B}}\sum_{w_{n}\in\mathcal{N}_{w}}\log(\text{similarity}(\phi_{s},w,w_{n}))$ (3)
The consistency loss aims to strengthen the alignment of anchor samples with their nearest neighbours, promoting cohesion by enhancing the similarity within neighbours.
We define an inconsistency loss to encourage dissimilarity between the anchor samples and their furthest neighbours in Equation (4). In this regard, we compute the pairwise similarity between the probability distributions of the anchor and furthest neighbour samples as well. Then, use the binary cross entropy loss to minimise the similarity to the furthest neighbours in the final loss function. While the mathematical form mirrors the consistency loss, its application diverges. Here, the similarity measure is used inversely; we seek to minimise this similarity, thereby driving a distinction between the anchor and its furthest neighbours to underscore cluster separation.
$\mathcal{L}_{inconsistency}(\phi_{s},\mathcal{B},\mathcal{F})=-\frac{1}{\lvert\mathcal{B}\rvert}\sum_{w\in\mathcal{B}}\sum_{w_{n}\in\mathcal{F}_{w}}\log(\text{similarity}(\phi_{s},w,w_{n}))$ (4)
To encourage class diversity and prevent overfitting, we use entropy loss on the distribution of anchor and neighbour samples across classes. Assuming the classes set is denoted as $\mathcal{C}=\{1,...,C\}$, and the probability of window $w_{i}$ being assigned to class $c$ is denoted as $\phi_{s}^{c}(w_{i})$:
$\mathcal{L}_{entropy}(\phi_{s},\mathcal{B},\mathcal{C})=\sum_{c\in\mathcal{C}}\hat{\phi_{s}^{c}}\log(\hat{\phi_{s}^{c}})\\ \text{ where }\hat{\phi_{s}^{c}}=\frac{1}{\lvert\mathcal{B}\rvert}\sum_{w_{i}\in\mathcal{B}}\phi_{s}^{c}(w_{i})$ (5)
The final objective in the Self-supervised Classification Stage is to minimise the total loss, calculated by the difference between the consistency and inconsistency losses, reduced by the entropy loss multiplied by a weight parameter, $\beta$:
$\mathcal{L}_{Self-supervised}(\phi_{s},\mathcal{B},\mathcal{N},\mathcal{F},\mathcal{C},\beta)=$ (6)
$\mathcal{L}_{consistency}(\phi_{s},\mathcal{B},\mathcal{N})-\mathcal{L}_{inconsistency}(\phi_{s},\mathcal{B},\mathcal{F})-\beta\cdot\mathcal{L}_{entropy}(\phi_{s},\mathcal{B},\mathcal{C})$
The goal of the loss function is to learn a representation that is highly discriminative, with nearest neighbours assigned to the same class as the anchor samples, and furthest neighbours assigned to a distinct class. By incorporating the semantically meaningful nearest and furthest neighbours, the model is able to produce more consistent and confident predictions, with the probability of a window being classified as one particular class is close to 1 or 0.
Overall, the loss function $\mathcal{L}_{Self-supervised}$ represents a critical component of our approach to time series anomaly detection, as it allows us to effectively learn a discriminative feature representation that can be utilised to distinguish between normal and anomalous windows.
3.4 CARLA’s Inference
Upon completing Part Two of our approach, we determine the class assignments for set $\mathcal{D}$ and majority class $C_{m}$, where $C_{m}=\arg\max_{C_{j}\in\mathcal{C}}(n(C_{j}))$ which comprises the class with the highest number of anchors (i.e. original windows). For every new window $w_{t}$ during inference, we calculate $\phi_{s}^{C_{m}}(w_{t})$, representing the probability of window $w_{t}$ being assigned to the majority class $C_{m}$. Using an end-to-end approach, we classify a given window $w_{t}$ as normal or anomalous based on whether it belongs to the majority class. Specifically, Equation (7) describes how we infer a label for a new time series window $w_{t}$. Additionally, we can employ Equation (8) to generate an anomaly score of $w_{t}$ for further analysis.
Anomaly label ($\text{Anomaly label }(w_{t}):\begin{cases}0,&\text{if }\forall c\in\mathcal{C},\phi_{s}^{C_{m}}(w_{t})\geq\phi_{s}^{c}(w_{t})\\ 1,&\text{otherwise}\end{cases}$ (7)
Anomaly score ($\text{Anomaly score }(w_{t}):1-\phi_{s}^{C_{m}}(w_{t})$ (8)
By assigning a window to a specific class, our model can determine the likelihood of a given window being normal or anomalous with greater precision.
Furthermore, the probability $\phi_{s}^{C_{m}}(w_{t})$ can be used to calculate an anomaly score, with lower values indicating a higher probability of anomalous behaviour. This score can be useful for further analysis and can aid in identifying specific characteristics of anomalous behaviour.
4 Experiments
The objective of this section is to thoroughly assess CARLA’s performance through experiments conducted on multiple benchmark datasets and to compare its results with multiple alternative methods. Section 4.1 provides an overview of the benchmark datasets used in the evaluation, highlighting their significance in assessing the effectiveness of our model. Section 4.2 delves into the benchmark methods employed for comparing the performance of different models. In section 4.3, we discuss the evaluation setup, including the hyper-parameters chosen for our approach. Moving forward, section 4.4 provides results for all benchmark methods based on the respective benchmark datasets, facilitating a comprehensive comparison. Additionally, we explore the behaviour of CARLA across diverse data configurations and variations in the ablation studies (section 4.5) and investigate its sensitivity to parameter changes in section 4.6. All evaluations were conducted on a system equipped with an A40 GPU, 13 CPUs, and 250 GB of RAM.
Table 1: Statistics of the benchmark datasets used.
Benchmark # datasets # dims Train size Test size Anomaly ratio
MSL 27 55 58,317 73,729 10.72%
SMAP 55 25 140,825 444,035 13.13%
SMD 28 38 708,405 708,420 4.16%
SWaT 1 51 496,800 449,919 12.33%
WADI 1 123 784,568 172,801 5.77%
Yahoo - A1 Benchmark 67 1 46,667 46,717 1.76%
KPI 29 1 1,048,576 2,918,847 1.87%
4.1 Benchmark Datasets
We evaluate the performance of the proposed model and make comparisons of the results across the seven most commonly used real benchmark datasets for TSAD. The datasets are summarised in Table 1, which provides key statistics for each dataset.
NASA DatasetsMars Science Laboratory (MSL) and Soil Moisture Active Passive (SMAP) [^1] https://www.kaggle.com/datasets/patrickfleith/nasa-anomaly-detection-dataset-smap-msl – (Hundman et al., 2018) are real-world datasets collected from NASA spacecraft. These datasets contain anomaly information derived from reports of incident anomalies for a spacecraft monitoring system. MSL and SMAP comprises of 27 and 55 datasets respectively and each equipped with a predefined train/test split, where, unlike other datasets, their training set are unlabeled.
Server Machine Dataset (SMD) [^2] https://github.com/NetManAIOps/OmniAnomaly/tree/master/ServerMachineDataset (Su et al., 2019) is gathered from 28 servers, incorporating 38 sensors, over a span of 10 days. During this period, normal data was observed within the initial 5 days, while anomalies were sporadically injected during the subsequent 5 days. The dataset is also equipped with a predefined train/test split, where the training data is unlabeled.
Secure water treatment (SWaT) [^3] https://itrust.sutd.edu.sg/testbeds/secure-water-treatment-swat/ (Mathur and Tippenhauer, 2016) is meticulously gathered within 11 days, utilising a water treatment platform that consisted of 51 sensors. During the initial 7 days, the dataset exclusively comprised normal data, accurately reflecting regular operational conditions. However, in the subsequent 4 days, a total of 41 anomalies were deliberately generated using a wide range of attack methods.
Water distribution testbed (WADI) [^4] https://itrust.sutd.edu.sg/testbeds/water-distribution-wadi/ (Ahmed et al., 2017) is acquired from a scaled-down urban water distribution system that included a combined total of 123 actuators and sensors. This data covers a span of 16 days. Notably, the final two days of the dataset specifically feature anomalies. In addition, the test dataset encompasses 15 distinct anomaly segments, providing valuable examples for anomaly detection and analysis purposes.
Yahoo [^5] https://webscope.sandbox.yahoo.com/catalog.php?datatype=s&did=70 dataset (Laptev et al., 2015) is a benchmark dataset that consists of hourly 367 sampled time series with anomaly points labels. The dataset includes four benchmarks, with the A1 benchmark containing ”real” production traffic data from Yahoo properties, and the other three are synthetic. For evaluation purposes, we focus on the A1 benchmark, which includes 67 univariate time series. However, some of these contain no anomalies in the test, hence we exclude them from the evaluation.
Key performance indicators (KPI) [^6] https://github.com/NetManAIOps/KPI-Anomaly-Detection contains time series data of service and machine key performance indicators (KPIs) from real scenarios of Internet companies. Service KPIs measure web service performance, while machine KPIs reflect machine health. Examples of service KPIs include page response time and page views, while machine KPIs include CPU and memory utilisation. The dataset includes 29 different KPI datasets, with separate training and test datasets provided.
4.2 Benchmark Methods
In this overview, provide a description of the ten prominent TSAD models that were used for comparison with CARLA, spanning various TSAD categories. For the semi-supervised approach, the LSTM-VAE (Park et al., 2018) is highlighted, which is a point-wise anomaly detection model. In the realm of unsupervised reconstruction-based models, Donut (Xu et al., 2018) stands out for UTS, while OmniAnomaly Su et al. (2019) and AnomalyTransformer (Xu et al., 2021a) are noted for MTS, alongside TranAD (Tuli et al., 2022) which caters to both UTS and MTS. The unsupervised forecasting-based model, THOC Shen et al. (2020), is recognized for its predictive capabilities. Within the self-supervised category, MTAD-GAT (Zhao et al., 2020) emerges as a hybrid model, with TS2Vec Yue et al. (2022) and DCdetector Yang et al. (2023) offering representation-based methodologies. Additionally, recent advancements are represented by TimesNet Wu et al. (2023) and DCdetector Yang et al. (2023), showcasing the ongoing evolution in TSAD models.
Donut  [^7] https://github.com/NetManAIOps/donut Xu et al. (2018) is a method for detection anomalies in UTS. It utilises a VAE for a reconstruction-based strategy. By learning the latent representation of time series data, it accurately identifies anomalies when data reconstruction significantly deviates from the original.
LSTM-VAE [^8] https://github.com/TimyadNyda/Variational-Lstm-Autoencoder Park et al. (2018) is a point-wise model that combines LSTM networks, which capture temporal data dependencies, with VAEs for generating high-dimensional data distributions and modelling complex data patterns. This integration allows for effective anomaly detection in MTS points by identifying deviations from normal behaviour, which is crucial in applications like robot-assisted feeding for ensuring safety and efficiency.
OmniAnomaly [^9] https://github.com/smallcowbaby/OmniAnomaly Su et al. (2019) is an unsupervised model that utilises a Variational Autoencoder (VAE) to model multivariate time series data. By analyzing the likelihood of a given window’s reconstruction, the model can detect anomalies.
MTAD-GAT [^10] https://github.com/mangushev/mtad-gat Zhao et al. (2020) enhances MTS anomaly detection by leveraging graph attention networks to understand complex dependencies across different time series. It employs a self-supervised framework that uses graph attention layers to learn relationships in both temporal and feature dimensions. This approach optimizes both forecasting-based and reconstruction-based models for anomaly detection.
THOC [^11] We utilised the authors’ shared implementation, as it is not publicly available. Shen et al. (2020) utilises a multi-layer dilated recurrent neural network (RNN) alongside skip connections in order to effectively handle contextual information. THOC utilise a temporal hierarchical one-class network for anomaly detection.
AnomalyTransformer [^12] https://github.com/thuml/Anomaly-Transformer Xu et al. (2021a) introduces an approach for detecting anomalies in MTS by leveraging the Transformer architecture, known for its effectiveness in capturing long-range dependencies. It employs association discrepancy to identify anomalies, focusing on the differences in relationships among data points in a time series.
TranAD [^13] https://github.com/imperial-qore/TranAD Tuli et al. (2022) leverages deep transformer networks for efficient anomaly detection in MTS and UTS. It employs attention-based sequence encoders for quick inference, leveraging focus score-based self-conditioning and adversarial training for robust feature extraction and stability.
TS2Vec [^14] https://github.com/yuezhihan/ts2vec Yue et al. (2022) is an unsupervised contrastive model that is capable of learning multiple contextual representations of MTS and UTS semantically at various levels. TS2Vec employs contrastive learning in a hierarchical way, which provides a contextual representation. They proposed a method to be employed for time series anomaly detection.
DCdetector [^15] https://github.com/DAMO-DI-ML/KDD2023-DCdetector Yang et al. (2023) introduces an approach for time series anomaly detection by employing dual attention mechanisms and contrastive representation learning. This method focuses on capturing intricate temporal relationships and distinguishing between normal and anomalous patterns in time series data.
TimesNet [^16] https://github.com/thuml/TimesNet Wu et al. (2023) proposes a framework for analyzing general time series data by modelling temporal 2D variations. This approach allows for capturing complex patterns and dependencies in time series. This method leverages advanced neural networks to enhance analysis accuracy and efficiency across various applications like time series anomaly detection.
Random Anomaly Score. The “Random Anomaly Score” model is designed to generate anomaly scores for time series windows based on a normal distribution ($\mathcal{N}(0,1)$) with mean 0 and standard deviation 1. This model has been specifically developed to showcase evaluation metrics and strategies in the field of TSAD.
4.3 Evaluation Setup
In our study, we evaluate ten TSAD models on benchmark datasets previously mentioned in section 4.1 using their best hyper-parameters as they stated to ensure a fair evaluation. We summarise the default hyper-parameters used in our implementation as follows: The CARLA model consists of a 3-layer ResNet architecture with three different kernel sizes [8, 5, 3] to capture temporal dependencies. The dimension of the representation is 128. We use the same hyper-parameters across all datasets to evaluate CARLA: window size = 200, number of classes = 10, number of nearest/furthest neighbours ($Q$) = 5, and coefficient of entropy = 5. For detailed information about all experiments involving the aforementioned hyper-parameter choices, please refer to the section 4.5. We run the Pretext Stage for 30 epochs and the Self-supervised Classification Stage for 100 epochs on all datasets.
It is important to note that we do not use Point Adjustment (PA) in our evaluation process. Despite its popularity, we do not use PA due to the findings in Kim et al. (2022) that its application leads to an overestimation of a TSAD model’s capability and can bias results in favour of methods that produce extreme anomaly scores. Thus, to ensure the accuracy and validity of our results, we relegate PA results to B and present here conventional F1 scores. The F1 score calculated without considering the PA is referred to as F1, while the adjusted predictions are denoted as $F1_{PA}$ in B.
Table 2: Precision (Prec), recall (Rec), F1 and AU-PR results for various models on multivariate time series datasets. The most optimal evaluation results are displayed in bold, while the second best ones are indicated by underline. Due to the single time series in both the SWaT and WADI datasets, the standard deviation of AU-PR is not available.
Model Metric MSL SMAP SMD SWaT WADI Avg Rank
LSTM-VAE Park et al. (2018) (2018) Prec 0.2723 0.2965 0.2045 0.9707 0.0596 5.5
Rec 0.8083 0.8303 0.5491 0.5769 1.0000
F1 0.4074 0.4370 0.2980 0.7237 0.1126
AU-PR 0.285$\pm$0.249 0.258$\pm$0.305 0.395$\pm$0.257 0.685 0.039
OmniAnomaly Su et al. (2019) (2019) Prec 0.1404 0.1967 0.3067 0.9068 0.1315 5.0
Rec 0.9085 0.9424 0.9126 0.6582 0.8675
F1 0.2432 0.3255 0.4591 0.7628 0.2284
AU-PR 0.149$\pm$0.182 0.115$\pm$0.129 0.365$\pm$0.202 0.713 0.120
MTAD-GAT Zhao et al. (2020) (2020) Prec 0.3559 0.3783 0.2473 0.1387 0.0706 4.2
Rec 0.7067 0.8239 0.5834 0.9585 0.5838
F1 0.4734 0.5186 0.3473 0.2423 0.1259
AU-PR 0.335$\pm$0.259 0.339$\pm$0.300 0.401$\pm$0.263 0.095 0.084
THOC Shen et al. (2020) (2020) Prec 0.1936 0.2039 0.0997 0.5453 0.1017 6.8
Rec 0.7718 0.8294 0.5307 0.7688 0.3507
F1 0.3095 0.3273 0.1679 0.6380 0.1577
AU-PR 0.239$\pm$0.273 0.195$\pm$0.262 0.107$\pm$0.126 0.537 0.103
AnomalyTransformer Xu et al. (2021a) (2021) Prec 0.2182 0.2669 0.2060 0.9711 0.0601 5.4
Rec 0.8231 0.8600 0.5822 0.5946 0.9604
F1 0.3449 0.4074 0.3043 0.7376 0.1130
AU-PR 0.236$\pm$0.237 0.264$\pm$0.315 0.273$\pm$0.232 0.681 0.040
TranAD Tuli et al. (2022) (2022) Prec 0.2957 0.3365 0.2649 0.1927 0.0597 6.2
Rec 0.7763 0.7881 0.5661 0.7965 1.0000
F1 0.4283 0.4716 0.3609 0.3103 0.1126
AU-PR 0.278$\pm$0.239 0.287$\pm$0.300 0.412$\pm$0.260 0.192 0.039
TS2Vec Yue et al. (2022) (2022) Prec 0.1832 0.2350 0.1033 0.1535 0.0653 7.7
Rec 0.8176 0.8826 0.5295 0.8742 0.7126
F1 0.2993 0.3712 0.1728 0.2611 0.1196
AU-PR 0.132$\pm$0.135 0.148$\pm$0.165 0.113$\pm$0.075 0.136 0.057
DCdetector Yang et al. (2023) (2023) Prec 0.1288 0.1606 0.0432 0.1214 0.1417 9.1
Rec 0.9578 0.9619 0.9967 0.9999 0.9684
F1 0.2270 0.2753 0.0828 0.2166 0.2472
AU-PR 0.129$\pm$0.144 0.124$\pm$0.153 0.043$\pm$0.036 0.126 0.121
TimesNet Wu et al. (2023) (2023) Prec 0.2257 0.2587 0.2450 0.1214 0.1334 6.1
Rec 0.8623 0.8994 0.5474 1.0000 0.1565
F1 0.3578 0.4019 0.3385 0.2166 0.1440
AU-PR 0.283$\pm$0.213 0.208$\pm$0.211 0.385$\pm$0.225 0.083 0.084
Random Prec 0.1746 0.1801 0.0952 0.1290 0.0662 8.4
Rec 0.9220 0.9508 0.9591 0.9997 0.9287
F1 0.2936 0.3028 0.1731 0.2166 0.1237
AU-PR 0.172$\pm$0.133 0.140$\pm$0.148 0.089$\pm$0.058 0.129 0.067
CARLA Prec 0.3891 0.3944 0.4276 0.9886 0.1850 1.6
Rec 0.7959 0.8040 0.6362 0.5673 0.7316
F1 0.5227 0.5292 0.5114 0.7209 0.2953
AU-PR 0.501$\pm$0.267 0.448$\pm$0.326 0.507$\pm$0.195 0.681 0.126
Table 3: Precision (Prec), recall (Rec), F1, and AU-PR results for various models on univariate time series datasets. The most optimal evaluation results are displayed in bold, while the second best ones are indicated by underline.
Model Metric Yahoo-A1 KPI Avg Rank
Donut Xu et al. (2018) (2018) Prec 0.3239 0.0675 5.0
Rec 0.9955 0.9340
F1 0.4888 0.1259
AU-PR 0.264$\pm$0.244 0.078$\pm$0.073
THOC Shen et al. (2020) (2020) Prec 0.1495 0.1511 5.5
Rec 0.8326 0.5116
F1 0.2534 0.2334
AU-PR 0.349$\pm$0.342 0.229$\pm$0.217
TranAD Tuli et al. (2022) (2022) Prec 0.4185 0.2235 2.0
Rec 0.8712 0.4016
F1 0.5654 0.2872
AU-PR 0.691$\pm$0.324 0.285$\pm$0.206
TS2Vec Yue et al. (2022) (2022) Prec 0.3929 0.1333 5.0
Rec 0.6305 0.4329
F1 0.4841 0.2038
AU-PR 0.491$\pm$0.352 0.221$\pm$0.156
DCdetector Yang et al. (2023) (2023) Prec 0.0598 0.0218 8.0
Rec 0.9434 0.8589
F1 0.1124 0.0425
AU-PR 0.041$\pm$0.059 0.018$\pm$0.017
TimesNet Wu et al. (2023) (2023) Prec 0.3808 0.2174 3.0
Rec 0.7883 0.2713
F1 0.5135 0.2414
AU-PR 0.671$\pm$0.307 0.237$\pm$0.148
Random Prec 0.2991 0.0657 6.5
Rec 0.9636 0.9488
F1 0.4565 0.1229
AU-PR 0.258$\pm$0.184 0.079$\pm$0.064
CARLA Prec 0.5747 0.1950 1.0
Rec 0.9755 0.7360
F1 0.7233 0.3083
AU-PR 0.645$\pm$0.352 0.299$\pm$0.245
4.4 Benchmark Comparison
The performance metrics employed to evaluate the effectiveness of all models consist of precision, recall, the traditional F1 score (F1 score without PA), FPR and the area under the precision-recall curve (AU-PR). Additionally, we computed the average ranks of the models based on their F1. For all methods, we used the precision-recall curve on the anomaly score for each time series in the benchmark datasets to find the best F1 score based on precision and recall for the target time series.
Since certain benchmark datasets such as MSL contain multiple time series datasets (as shown in Table 1), we cannot merge or combine these time series due to the absence of timestamp information. Additionally, calculating the F1 score for the entire dataset by averaging individual scores is not appropriate. In terms of precision and recall, the F1 score represents the harmonic mean, which makes it a non-additive metric. To address this, we get the confusion matrix for each time series, i.e., we calculate the number of true positives (TP), false positives (FP), true negatives (TN), and false negatives (FN) for each dataset of a benchmark such as MSL. Then, we sum up the TP, FP, TN, and FN values from all the confusion matrices to get an overall confusion matrix of the entire benchmark dataset. After that, we use the overall confusion matrix to calculate the overall precision, recall, and F1 score. This way, we ensure that the F1 score is calculated correctly for the entire dataset rather than being skewed by averaging individual F1 scores.
Furthermore, for datasets with more than one time series, we report the average and standard deviation of AU-PR across all time series. AU-PR is advantageous in imbalanced datasets because it is less sensitive to the distribution of classes Saito and Rehmsmeier (2015). It considers the trade-off between precision and recall across all possible decision thresholds, making it more robust in scenarios where the number of instances in classes are imbalanced. By focusing on the positive class (i.e. anomalous) and its predictions, AU-PR offers a more precise evaluation of a model’s ability to identify and prioritise the minority class correctly.
Tables 2 and 3 demonstrate the performance comparison between CARLA and all benchmark methods for multivariate and univariate datasets, respectively. These two tables show our model outperforms other models with higher F1 and AU-PR scores across all datasets (except SWaT for F1 and AU-PR and Yahoo-A1 for AU-PR), in which CARLA is the third best. This shows the strength of CARLA in generalising normal patterns due to its high precision. CARLA’s precision is the highest across all seven datasets. At the same time, CARLA has high enough recall, and as a result, our CARLA’s F1 is the highest across all datasets except SWaT. While the recall is higher in other benchmarks, such as OmniAnomaly, it is with the expense of very low precision (with median precision $<19.67\%$ across all datasets). This means that a good balance is achieved between precision and recall by CARLA. This is also shown in our consistently high AU-PR compared to other models and indicates that it is proficient in accurately identifying anomalous instances with higher recall while minimising false alarms with higher precision on imbalanced datasets where normal instances are predominant. Finally, CARLA achieved the lowest average rank (best rank), based on F1, in both multivariate and univariate benchmarks (see last column in Tables 2 and 3).
Figure 4: Models’ performance comparison on multivariate time series datasets regarding their FPR and AU-PR. Models in the optimal quadrant (low FPR, high AU-PR) demonstrate superior anomaly detection with minimal false alarms.
Figure 5: Models’ performance comparison on univariate time series datasets regarding their FPR and AU-PR. Models in the optimal quadrant (low FPR, high AU-PR) demonstrate superior anomaly detection with minimal false alarms.
We now evaluate the performance of all models in terms of FPR. We also incorporate AU-PR in this comparison, as some models can achieve a low FPR at the expense of lowering AU-PR. In the scatter plots provided in Figure 4 we show FPR on the x-axis and AU-PR on the y-axis. The ideal point on this scatter plot is in the leftmost top corner. CARLA consistently shows a lower FPR compared to other models across various multivariate time series datasets. This is particularly noteworthy in the context of anomaly detection, where the cost of false positives can be substantial, leading to wasted resources and potential overlook of true anomalies. CARLA’s position in the plots are close to leftmost top corner (in the optimal quadrant). This indicates fewer false alarms while maintaining competitive AU-PR scores, which measures the precision and recall balance of anomaly detection.
For instance, on the MSL dataset, CARLA is among the models with the lowest FPR, and its AU-PR score is in the upper tier, showcasing its ability to identify true anomalies accurately. On the SMAP and SMD datasets, CARLA maintains a significantly lower FPR than most models, suggesting its robustness to varying data conditions. In the SWaT dataset, CARLA excels with the lowest FPR and the third highest AU-PR, marking it suitable for precision-critical applications. Even on WADI, while CARLA’s FPR is slightly higher, it remains in the optimal quadrant, reflecting a strong FPR and AU-PR balance vital for maintaining operational continuity and effective anomaly detection.
Adding to CARLA’s impressive performance on multivariate datasets, its capabilities in univariate time series analysis, as shown in Figure 5, further highlighting its robustness. On the Yahoo-A1 dataset, while CARLA does not achieve the highest AU-PR, it still demonstrates a commendable balance of a low FPR and a high AU-PR, positioning it as the third-best model in this regard. Its place in the upper left quadrant indicates a strong ability to correctly identify anomalies with a minimal rate of false alarms. On the KPI dataset, CARLA maintains a competitive edge, presenting a low FPR that surpasses most other models. While the AU-PR is the highest, CARLA’s ability to limit false positives is a notable strength.
Overall, the consistent performance of CARLA across multivariate and univariate datasets and its proficiency in dealing with a variety of data types, as shown by its low FPR, highlights its strength as a reliable model for anomaly detection in MTS and UTS data. Its ability to minimise false positives without significantly sacrificing true positive detection makes it an excellent choice for scenarios where high-confidence alerts are vital.
4.5 Ablation Study
Additionally, for a deeper understanding of the individual contributions of the stages and different components in our proposed model, we conduct an ablation study. Our analysis focused on the following aspects: (i) Effectiveness of CARLA’s two stages, (ii) Positive pair selection strategy (iii) Effectiveness of Different Anomaly Types (iv) Effectiveness of the loss components.
4.5.1 Effectiveness of CARLA’s Stages
In this section, we delve into an in-depth evaluation of the stages of our proposed model, utilising the M-6 time series derived from the MSL dataset. Our primary objective is to develop a comprehensive grasp of the patterns inherent in the features extracted and learned by our model throughout its stages. For this purpose, we employ t-distributed Stochastic Neighbor Embedding (t-SNE) to visualise the output of the Pretext Stage and the Self-supervised Classification stage. These are graphically represented in Figures 5(a) and 5(b) correspondingly. From these graphical representations, it is palpable that the second stage significantly enhances the discrimination capacity between normal and anomalous samples. It is important to highlight that the first stage reveals some semblance of anomalous samples to normal ones, which might be due to the existence of anomalies close to normal boundaries. However, the Self-supervised Classification stage efficiently counteracts this ambiguity by effectively segregating these instances, thereby simplifying subsequent classification tasks.
Figure 6: Comparative analysis of the model’s stages utilising t-SNE and anomaly score distributions on M-6 dataset in MSL. (a) t-SNE after the Pretext stage, indicating the initial representation of normal and anomalous samples. (b) t-SNE after the Self-supervised Classification stage, demonstrating enhanced segregation of normal and anomalous instances. (c) Anomaly score distribution in the Pretext Stage based on Euclidean distances between test and training samples. (d) The distribution of anomaly scores following the application of the Self-supervised Classification Stage, evidencing an improved representation.
Further, we assessed the efficacy of the Self-supervised Classification Stage in refining the representation generated by the Pretext Stage. This was accomplished by juxtaposing the anomaly scores derived from the output of each stage. In Figure 5(c), the distribution of anomaly scores for the test windows at the Pretext Stage is displayed, computed using the Euclidean distances of the test samples relative to the original time series windows in the training set. Conversely, Figure 5(d) depicts the subsequent alteration in the distribution of anomaly scores following the application of the Self-supervised Classification Stage.
In the Pretext Stage, the anomaly score is computed as the minimum distance between the test sample and all original training samples in the representation space. This approach aims to identify the closest match among the training samples for the given test sample. A smaller distance implies a greater probability that the test sample belongs to the normal data distribution.
In the Self-supervised Classification Stage, we use the inference step in Section 3.3 of the article. Equation 8 provides further details on the computation of the anomaly label.
As we can observe from the figures 5(a) and 5(b), the self-supervised classification Stage has resulted in a significant improvement in the separation of normal and anomalous windows. The distribution of anomaly scores in Figure 5(d) is more clearly separated than in Figure 5(c). This indicates that the second stage has produced more consistent representations. Furthermore, the anomaly scores in Figure 5(d) are relatively closer to 0 or 1, which indicates an improvement in discrimination between normal and anomalous windows in the second stage of CARLA.
4.5.2 Positive Pair Selection
To evaluate the effectiveness of the positive pair selection method in CARLA, we compared two approaches in training: random temporal neighbour from $y$ temporally closest window samples and weak augmentation with noise (add normal noise with sigma 0.01 to a window). We used MSL benchmark and used identical configuration and hyper-parameters for both approaches and evaluated their performance shown in Table 4.
Table 4: Positive pair selection results on MSL dataset.
Pos Pair Selection Prec Rec F1 AU-PR
CARLA-noise 0.3086 0.7935 0.4433 0.3983
CARLA-temporal 0.3891 0.7959 0.5227 0.5009
Our experiment demonstrated that selecting positive pairs using a random temporal neighbour is a more effective approach than weak augmentation with noise. Since anomalies occur rarely, choosing a positive pair that is in the target window’s temporal proximity is a better strategy in our model.
4.5.3 Effectiveness of Different Anomaly Types
Based on the results of the ablation study on the MSL dataset, CARLA’s performance is evaluated by systematically removing different types of anomalies during the Pretext Stage. The evaluation metrics used are precision, recall, F1 score, and AU-PR. The results are sorted based on F1 from the least significant to the most significant types of anomalies in Table 5. This result represents the overall performance of the CARLA model when using all types of anomalies during the anomaly injection process. It serves as a baseline for comparison with the subsequent results.
Table 5: Effectiveness of different anomaly types on MSL dataset.
Anomaly Type Prec Rec F1 AU-PR
All types 0.3891 0.7959 0.5227 0.5009
w/o trend 0.3185 0.7849 0.4513 0.4972
w/o contextual 0.3842 0.7935 0.5177 0.4683
w/o shapelet 0.3564 0.8370 0.4999 0.4328
w/o global 0.3383 0.8404 0.4824 0.4340
w/o seasonal 0.2400 0.8428 0.3737 0.3541
In the experiments, anomalies were injected into the training data on a per-window basis. For each window in the multivariate time series, a random number of dimensions was selected, ranging from 1 to $\lceil Dim/10\rceil$ of the total dimensions. These selected dimensions were injected with anomalies, starting from the same point across all the chosen dimensions. The injected anomaly portion for each dimension varied from 1 data point to 90% of the window length. This approach ensured a diverse and controlled injection of anomalies within the training data (For more detail, see Algorithm 1)
Eliminating trend anomalies has a significant negative impact on CARLA’s precision. However, the recall remains relatively high, indicating that trend anomalies are important for maintaining a higher precision level. Similar to the previous case, removing contextual anomalies leads to a slight decrease in precision but an improvement in recall. The exclusion of shapelet anomalies leads to a moderate decrease in precision, while the recall remains relatively high. This suggests that shapelet anomalies contribute to the model’s precision but are not as crucial for capturing anomalies in general.
By removing global anomalies, CARLA’s precision slightly decreases, indicating that it becomes less accurate in identifying true anomalies. However, the recall improves, suggesting that the model becomes more sensitive in detecting anomalies overall. Removing seasonal anomalies has a noticeable effect on the model’s precision, which drops significantly. The consistently high recall suggests the model effectively captures many anomalies without relying on seasonal patterns.
Overall, the analysis of the results suggests that global and seasonal anomalies play a relatively more significant role in the CARLA model’s performance, while contextual, shapelet, and trend anomalies have varying degrees of impact. These findings provide insights into the model’s sensitivity to different types of anomalies and can guide further improvements in anomaly injection strategies within CARLA’s framework.
4.5.4 Effectiveness of Loss Components
To assess the effectiveness of two loss components, namely $\mathcal{L}_{inconsistency}$ and $\mathcal{L}_{entropy}$, in the total loss function for self-supervised classification on the MSL dataset, we conduct experiments using identical architecture and hyper-parameters. The evaluation metrics employed to measure the performance are precision, recall, F1, and AU-PR, as presented in Table 6.
Table 6: Effectiveness of $\mathcal{L}_{inconsistency}$ and $\mathcal{L}_{entropy}$ on MSL dataset.
$\mathcal{L}_{inconsistency}$ $\mathcal{L}_{entropy}$ Prec Rec F1 AU-PR
0.3092 0.7155 0.4318 0.3984
0.3453 0.8112 0.4846 0.4424
0.3219 0.7534 0.4511 0.4175
0.3891 0.7959 0.5227 0.5009
In cases where only one loss component is used, the model’s performance is relatively lower across all metrics, indicating that without incorporating these loss components, the model struggles to detect anomalies in the MSL dataset effectively.
Where both $\mathcal{L}_{inconsistency}$ and $\mathcal{L}_{entropy}$ loss components are included in the total loss function, the model achieves the best performance among all scenarios, surpassing the other combinations. The precision, recall, F1 score, and AU-PR are the highest in this case, indicating that the combination of both loss components significantly improves the anomaly detection capability of the model.
4.6 Parameter Sensitivity
We also examine the sensitivity of CARLA’s parameters. We investigate and analyze the effect of the window size, number of classes, number of NN/FN, and coefficient of entropy.
4.6.1 Effect of Window Size
As a hyper-parameter in time series analysis, window size holds considerable importance. Table 7 and Figure 7 present the results of exploring the impact of window size on three datasets: MSL, SMD, and Yahoo.
Figure 7: (a) F1 Score and (b) AU-PR for different Window sizes across MSL, SMD, and Yahoo datasets
Table 7: Exploring the effect of window size on MSL, SMD, and Yahoo datasets.
MSL SMD Yahoo
WS
Prec
Rec
F1
AUPR
Prec
Rec
F1
AUPR
Prec
Rec
F1
AUPR
50
0.2755
0.7702
0.4058
0.3173
0.4569
0.4885
0.4722
0.4636
0.3046
0.8800
0.4526
0.5533
100
0.3233
0.7666
0.4548
0.4155
0.2584
0.5938
0.3598
0.4136
0.3709
0.9713
0.5368
0.5331
150
0.3606
0.7823
0.4936
0.4595
0.3599
0.6592
0.4655
0.4661
0.5720
0.9578
0.7162
0.6685
200
0.3891
0.7959
0.5227
0.5009
0.4276
0.6362
0.5114
0.5070
0.5747
0.9755
0.7233
0.6450
250
0.4380
0.8037
0.5412
0.5188
0.4224
0.6596
0.5150
0.4850
0.7129
0.9781
0.8247
0.7164
Based on the analysis of both the F1 score and AU-PR, it can be concluded that window size 200 consistently outperforms other window sizes on the MSL, SMD, and Yahoo datasets overall. This window size strikes a balance between precision and recall, effectively capturing anomalies while maintaining a high discrimination ability. Therefore, window size 200 is selected as the best choice for conducting time series analysis in all experiments. Additionally, Figure 7 shows that our model’s performance is stable in window sizes between 150 and 250.
Figure 8: F1 Score and AU-PR for (a) different number of classes, (b) different numbers of neighbours in $\mathcal{N}$$\mathcal{F}$ and (c) different coefficient of entropy, on MSL dataset.
4.6.2 Effect of Number of Classes
The provided results showcase the performance of the model with varying numbers of classes on the MSL dataset. Based on the illustration of both the F1 score and AU-PR in Figure 7(a), the model performs relatively well across different numbers of classes. While CARLA’s performance is pretty stable across all number of classes denoted on the x-axis of the plot in Figure 7(a), it can be concluded that using 10 classes as the number of class parameters yields the highest performance. It strikes a balance between capturing anomalies and minimising false positives while effectively discriminating between anomalies and normal data points.
The results suggest that increasing the number of classes beyond 10 does not significantly improve the model’s performance. When the number of classes increases, normal representations divide into the different classes and, in CARLA, are detected as anomalies (lower probability of belonging to the major class). However, using 2 classes leads to a lower performance compared to the selected parameter of 10 classes, implying that a more fine-grained classification with 10 classes provides the assumption that various anomalies are spread in the representation space.
4.6.3 Effect of Number of neighbours in $\mathcal{N}$ and $\mathcal{F}$
The provided results in Figure 7(b) display the evaluation metrics for different numbers of nearest neighbours ($\mathcal{N}$) and furthest neighbours ($\mathcal{F}$) noted as $Q$.
Based on the analysis of both the F1 score and AU-PR, we can see that while CARLA’s performance is pretty stable across the different numbers of neighbours in $\mathcal{N}$ and $\mathcal{F}$ denoted on the x-axis of the plot in Figure 7(b), it can be concluded that employing 5 nearest/furthest neighbours outperforms the other options. This number of parameters strikes a favourable balance between precision and recall, allowing for effective anomaly detection while maintaining a high discrimination ability. Therefore, selecting 5 for the number of neighbours is deemed the optimal choice for this experiment, as it consistently achieves higher F1 scores and AU-PR values.
4.6.4 Effect of Coefficient of Entropy
In this experiment, we explore the impact of the entropy coefficient. Figure 7(c) shows the results of evaluating the coefficient of the entropy component in the loss function on the MSL dataset. The coefficients tested are 0, 2, 5, and 10.
Based on the analysis of both the F1 score and AU-PR, a coefficient value of 5 emerges as the most effective choice for the entropy component in the loss function. It achieves the highest scores for both metrics, indicating better precision-recall balance and accurate ranking of anomalies. However, it is worth noting that coefficients of 2 and 10 also demonstrate competitive performance, slightly lower than the value of 5. The coefficient of 0 significantly reduces the model’s performance, further emphasising the importance of incorporating the entropy component for improved anomaly detection performance.
Overall, the analysis indicates that coefficients of 2, 5, and 10 are viable for the entropy component in the loss function, but a coefficient of 5 is chosen as optimal due to its superior F1 score and AU-PR.
5 Conclusion
In conclusion, our innovative end-to-end self-supervised framework, CARLA, utilises contrastive representation learning with anomaly injection to generate robust representations for time series windows and classify anomalous windows. The use of semantically meaningful nearest and furthest neighbours as a prior allows us to capture underlying patterns in time series data and learn a representation that is well-aligned with these patterns, thereby enhancing the accuracy of anomaly detection. Additionally, the experimental evaluation shows promising results in detecting anomalies in the seven most commonly used real-world time series datasets, demonstrating the effectiveness of CARLA in this domain.
This research has demonstrated the potential of contrastive learning combined with synthetic anomaly injection to overcome the limitations of the lack of labelled data in time series anomaly detection. To enhance our model’s performance, future research will focus on refining positive and negative sample selection strategies. Moreover, we believe that there is much scope to improve performance by broadening the forms of injected anomalies used in our framework. We further hypothesise that varying the severity of injected anomalies may improve the model’s ability to detect anomalies in representation space. We commend the investigation of these intriguing prospects in future research.
Additionally, considering that some datasets with a significant number of abnormal samples, resulting in more anomalous windows during the Pretext Stage, it becomes critical to revisit the impact of selecting positives and negatives for abnormal anchor samples under unsupervised settings. This reconsideration is crucial for enhancing the model’s effectiveness in time series anomaly detection.
In conclusion, this paper’s results, based on appropriate metrics, contribute significantly to further investigations and advancements in time series anomaly detection, particularly on datasets like WADI and KPI. By employing the correct evaluation metrics, such as the conventional F1 score without point adjustment and AU-PR, this research provides a solid foundation for assessing the performance of future models in this field. The findings presented herein enable researchers to benchmark their models effectively and compare their results against established standards. Reliable evaluation metrics and validated results serve as valuable resources, driving progress and innovation in time series anomaly detection research.
Appendix A List of Symbols and Definitions
Table 8 demonstrates symbols used in the paper.
Table 8: List of Symbols and Definitions
Symbol Definition
$\mathcal{D}$
Sequential time series windows where $\mathcal{D}=\{w_{1},w_{2},..,w_{m}\}$
$w_{i}$
$i^{th}$ window in $\mathcal{D}$ where $w_{i}=\{x_{1},x_{2},..,x_{WS}\}$
$WS$
Time series window size
$Dim$
Number of dimensions in time series $\mathcal{D}$
$m$
Number of windows
$Q$
Number of nearest/furthest neighbours
$a$
Anchor sample (i.e. the original time series window)
$y$
Number of temporally proximate windows
$p$
Positive sample (i.e. random window selected from $y$ temporally proximate previous windows)
$n$
Negative sample (i.e. anomaly injected window)
$(a,p,n)$
Triplet in Pretext Stage
$\mathcal{T}$
All triplets set in Pretext Stage
$\alpha$
Margin in pretext loss
$\phi_{p}$
Pretext trained model
$\mathcal{B}$
All neighbours set
$\mathcal{N}$
$Q$-Nearest neighbours set
$\mathcal{F}$
$Q$-Furthest neighbours set
$\phi_{s}$
Self-supervised trained model
$C_{m}$
Majority class label
$\beta$
Entropy loss weight
Appendix B Point Adjustment (PA)
A notable issue with many existing Time Series Anomaly Detection (TSAD) methods is their reliance on a specific evaluation approach called point adjustment (PA) strategy (Xu et al., 2018) and utilised in TSAD literature such as Su et al. (2019). PA operates under the following principle: if a given segment is detected as anomalous, all previous segments in that anomaly subsequence are considered as anomalies (see Figure 9). PA was introduced under the assumption that a solitary alert during an anomaly phase is adequate to recover the system. It has evolved into a fundamental stage in the assessment of TSAD, with certain studies exclusively documenting $F1_{PA}$ without the conventional F1 score Kim et al. (2022). A higher $F1_{PA}$ is often interpreted as indicating better detection capability.
Figure 9: Impact of point adjustment (PA) on anomaly detection in time series. The figure represents a toy example using a sequence of 12 samples. Presents the anomaly scores for each sample within the sequence. The second row depicts the outcomes of anomaly detection using a threshold of 0.5 applied to all samples’ anomaly scores. The third row provides the actual labels (ground truth) for the samples within the sequence. The fourth and final row presents the adjusted anomaly detection results using point adjustment with the ground truth. Evaluation metrics are calculated for both unadjusted and adjusted anomaly detection.
In Figure 9, for the unadjusted detection, we have True Positives (TP): 2, True Negatives (TN): 3, False Positives (FP): 1, False Negatives (FN): 6, yielding a Recall of 0.25, Precision of 0.67, and F1 score of 0.36. In contrast, the adjusted detection (in bold for emphasis) resulted in TP: 7, TN: 3, FP: 1, and FN: 1, giving a significantly improved Recall of 0.875, Precision of 0.875, and F1 score of 0.875. This illustrates the substantial difference in evaluation results due to the application of PA in the anomaly detection process. It is evident that the point adjustment greatly enhances the performance of anomaly detection.
Table 9: Precision ($Prec_{PA}$), recall ($Rec_{PA}$), and $F1_{PA}$ results with point adjustment (PA) for various models on multivariate time series datasets.
Model Metric MSL SMAP SMD SWaT WADI
LSTM-VAE $Prec_{PA}$ 0.8049 0.6755 0.7765 0.9745 0.0639
$Rec_{PA}$ 0.8394 0.9358 0.8740 0.7036 1.0000
$F1_{PA}$ 0.8218 0.7848 0.8224 0.8172 0.1202
OmniAnomaly $Prec_{PA}$ 0.8241 0.6931 0.8193 0.9861 0.2685
$Rec_{PA}$ 0.9859 0.9754 0.9531 0.7864 0.9686
$F1_{PA}$ 0.8978 0.8104 0.8811 0.8750 0.4204
MTAD-GAT $Prec_{PA}$ 0.4305 0.7561 0.4106 0.9720 0.2553
$Rec_{PA}$ 0.9953 0.9957 0.9370 0.6857 0.9941
$F1_{PA}$ 0.6011 0.8595 0.5710 0.8052 0.4062
THOC $Prec_{PA}$ 0.8291 0.5741 0.3819 0.7854 0.3871
$Rec_{PA}$ 0.9651 0.9324 0.8946 0.9774 0.7638
$F1_{PA}$ 0.8919 0.7106 0.5353 0.8709 0.5138
AnomalyTransformer $Prec_{PA}$ 0.7201 0.6481 0.8058 0.9087 0.0642
$Rec_{PA}$ 0.9832 0.9258 0.8482 0.7735 1.0000
$F1_{PA}$ 0.8313 0.7625 0.8265 0.8357 0.1207
TranAD $Prec_{PA}$ 0.8911 0.8644 0.8043 0.5207 0.0639
$Rec_{PA}$ 0.9953 0.9976 0.9374 0.7779 1.0000
$F1_{PA}$ 0.9403 0.8906 0.9168 0.6238 0.1202
TS2Vec $Prec_{PA}$ 0.2152 0.2605 0.1800 0.1718 0.0893
$Rec_{PA}$ 1.0000 1.0000 0.9950 1.0000 1.0000
$F1_{PA}$ 0.3542 0.4133 0.3050 0.2932 0.1639
DCdetector $Prec_{PA}$ 0.4972 0.5648 0.6359 0.6998 0.2257
$Rec_{PA}$ 1.0000 0.9905 0.8110 0.9125 0.9263
$F1_{PA}$ 0.6642 0.7194 0.7128 0.7921 0.3630
TimesNet $Prec_{PA}$ 0.7547 0.6015 0.7962 0.8421 0.3742
$Rec_{PA}$ 0.9765 0.9822 0.9265 1.000 0.9127
$F1_{PA}$ 0.8514 0.7461 0.8560 0.9174 0.5308
Random $Prec_{PA}$ 0.1746 0.9740 0.8670 0.9398 0.8627
$Rec_{PA}$ 1.0000 1.0000 0.9686 0.9331 0.9732
$F1_{PA}$ 0.9840 0.9869 0.9150 0.9364 0.9146
CARLA $Prec_{PA}$ 0.8545 0.7342 0.6757 0.9891 0.1971
$Rec_{PA}$ 0.9650 0.9935 0.8465 0.7132 0.7503
$F1_{PA}$ 0.9064 0.8444 0.7515 0.8288 0.3122
Table 10: Precision ($Prec_{PA}$), recall ($Rec_{PA}$), and $F1_{PA}$ results with point adjustment (PA) for various models on univariate time series datasets.
Model Metric Yahoo-A1 KPI
Donut $Prec_{PA}$ 0.5252 0.3928
$Rec_{PA}$ 0.9280 0.5797
$F1_{PA}$ 0.6883 0.4683
THOC $Prec_{PA}$ 0.3138 0.3703
$Rec_{PA}$ 0.9723 0.9154
$F1_{PA}$ 0.4745 0.5273
TranAD $Prec_{PA}$ 0.5722 0.7012
$Rec_{PA}$ 0.9858 0.7528
$F1_{PA}$ 0.7241 0.7261
TS2Vec $Prec_{PA}$ 0.4993 0.4514
$Rec_{PA}$ 0.9717 0.9058
$F1_{PA}$ 0.6597 0.6025
DCdetector $Prec_{PA}$ 0.0750 0.2758
$Rec_{PA}$ 0.7653 0.8380
$F1_{PA}$ 0.1366 0.4151
TimesNet $Prec_{PA}$ 0.6232 0.5185
$Rec_{PA}$ 0.8865 0.6133
$F1_{PA}$ 0.7319 0.5619
Random $Prec_{PA}$ 0.9020 0.6540
$Rec_{PA}$ 0.9911 0.8986
$F1_{PA}$ 0.9445 0.7570
CARLA $Prec_{PA}$ 0.7475 0.3980
$Rec_{PA}$ 0.9984 0.8933
$F1_{PA}$ 0.8549 0.5507
However, PA carries a notable peril of inflating the performance of the model. In most of the TSAD models, an anomaly score is produced to indicate the degree of abnormality, and the model detects a given point as an anomaly when the score surpasses a predefined threshold. Incorporating PA, the predictions based on the model that randomly generated anomaly scores uniformly outperform all state-of-the-art models, as depicted in Tables 9 and 10. It becomes challenging to assert that a model with a superior $F1_{PA}$ performs better than a model with a lower $F1_{PA}$ when random anomaly scores have the potential to result in a high $F1_{PA}$ comparable to that of a proficient detection model. Experimental results in Tables 9 and 10 demonstrate that the random anomaly score model can outperform all state-of-the-art models, highlighting the limitations of relying solely on $F1_{PA}$ for model evaluation.

### v4

@@
 Additionally, it leverages the information about representations’ neighbours through a self-supervised approach to classify windows based on their nearest/furthest neighbours to further enhance the performance of anomaly detection.
-In extensive tests on seven major real-world time series anomaly detection datasets, CARLA shows superior performance (F1 and AU-PR) over state-of-the-art self-supervised, semi-supervised, and unsupervised TSAD methods for univariate time series and multivariate time series.
-Our research highlights the immense potential of contrastive representation learning in advancing the time series anomaly detection field, thus paving the way for novel applications and in-depth exploration in the domain.
+In extensive tests on seven major real-world TSAD datasets, CARLA shows superior performance (F1 and AU-PR) over state-of-the-art self-supervised, semi-supervised, and unsupervised TSAD methods for univariate time series and multivariate time series.
+Our research highlights the immense potential of contrastive representation learning in advancing the TSAD field, thus paving the way for novel applications and in-depth exploration.
 keywords:
 Anomaly Detection , Time Series , Deep Learning , Contrastive Learning , Representation Learning , Self-supervised learning
+journal: Elsevier
 […]
@@
 outliers) to ensure system reliability.
-Several machine learning algorithms are well-suited for detecting these outliers Domingues et al.
-(2018).
+Several machine learning algorithms are well-suited for detecting these outliers [7].
 In time series data, anomalies can result from different factors, including equipment failure, sensor malfunction, human error, and human intervention.
@@
 Time series anomaly detection (TSAD) has been the subject of decades of intensive research, with numerous approaches proposed to address the challenge of spotting rare and unexpected events in complex and noisy data.
-Statistical methods have been developed to monitor and identify abnormal behaviour Domingues et al.
-(2018).
-Recent advancements in deep learning techniques have effectively tackled various anomaly detection problems Pang et al.
-(2021).
-Specifically, for time series with complex nonlinear temporal dynamics, deep learning methods have demonstrated remarkable performance Schmidl et al.
-(2022).
+Statistical methods have been developed to monitor and identify abnormal behaviour [7].
+Recent advancements in deep learning techniques have effectively tackled various anomaly detection problems [22].
+Specifically, for time series with complex nonlinear temporal dynamics, deep learning methods have demonstrated remarkable performance [25].
 Most of these models focus on learning the normal behaviour of data from an unlabelled dataset and, therefore, can potentially predict samples that deviate from the normal behaviour as anomalies.
@@
 Models that cannot discriminate between normal and anomaly classes may predict normal samples as anomalous, leading to high false positives.
-Figure 0(a) shows an example of a histogram of anomaly scores of an existing unsupervised TSAD method called THOC Shen et al.
-(2020) on a benchmark dataset (M-6 test set from the MSL dataset).
+Figure 1(a) shows an example of a histogram of anomaly scores of an existing unsupervised TSAD method called THOC [26] on a benchmark dataset (M-6 test set from the MSL dataset).
 Frequent false positives often occur when normal data are assigned high anomaly scores.
 In instances where the representation of a segment of the normal data closely resembles that of anomalous data, there can be an elevated anomaly score attributed to these normal samples, consequently leading to an increase in false positives.
-This issue can be seen in right side of Figure 0(a) where many normal samples are assigned high anomaly scores.
+This issue can be seen on the right side of Figure 1(a), where many normal samples are assigned high anomaly scores.
 An alternative approach is to leverage self-supervised contrastive representation learning.
 Contrastive learning entails training a model to differentiate between pairs of similar and dissimilar samples in a dataset.
-Contrastive representation learning has been successful in image Chen et al.
-(2020) and natural language processing Logeswaran and Lee (2018), and its potential in TSAD has been explored in recent years.
-Contrastive loss functions can increase the distance between normal and their corresponding anomalous samples while decreasing the distance between normal samples.
+Contrastive representation learning has been successful in image [4] and natural language processing [18], and its potential in TSAD has been explored in recent years.
+In the context of an anomaly detection task where the majority of the data is normal, contrastive loss functions can increase the distance between normal samples and their corresponding negative (anomalous) samples.
+Simultaneously, they decrease the distance between normal samples and their corresponding positive (normal) samples.
 By doing so, the distinction between normal and anomalous data is made clearer and more pronounced, which helps to identify a more accurate boundary for the normal samples.
@@
 We argue that these assumptions carry the risk that augmentation of time series can transform them to negative samples, and a temporally distant window can represent a positive sample, leading to ineffective anomaly detection performance.
-Figure 0(b) shows anomaly scores of a contrastive learning TSAD method called TS2Vec Yue et al.
-(2022).
+Figure 1(b) shows anomaly scores of a contrastive learning TSAD method called TS2Vec [38].
 As demonstrated, the anomaly scores of both normal and anomalous data are blended together, resulting in the inefficacy of anomaly detection.
 In fact, in situations where normal and anomalous data are intermingled within the representation space, achieving an enhanced anomaly detection rate—as measured by metrics such as the F1 score or the area under the precision-recall curve (AU-PR)—is accompanied by an increase in the false positive rate.
-Figure 1: Histogram of anomaly scores produced by (a) THOC model Shen et al.
-(2020) and (b) TS2Vec model Yue et al.
-(2022).
-Histograms illustrating the distribution of anomaly scores for the THOC and TS2Vec models using data from entity M-6 of the MSL dataset Hundman et al.
-(2018).
+Figure 1: Histograms of the distribution of anomaly scores produced by (1(a)) THOC [26], (1(b)) and TS2Vec [38] and (1(c)) CARLA models using M-6 dataset of the MSL benchmark [12].
 We propose a novel two-stage framework called CARLA, designed specifically to enhance time series anomaly detection.
-Our novel approach addresses the lack of labelled data through a contrastive approach which leverages existing generic knowledge about different types of time series anomalies in the first stage (pretext).
+Our novel approach addresses the lack of labelled data through a contrastive approach, which leverages existing generic knowledge about different types of time series anomalies [5] in the first stage (pretext).
 We inject various types of anomalies, which facilitates the learning representations of normal behaviour.
@@
     We propose a novel contrastive representation learning model to detect anomalies in time series, which delivers top-tier outcomes across a range of real-world benchmark datasets, encompassing both univariate time series (UTS) and multivariate time series (MTS).
-Addressing the challenge of lack of labelled data, our model learns to effectively discriminative normal patterns from anomalous in the feature representation space.
+Addressing the challenge of lack of labelled data, our model learns to effectively discriminate normal patterns from anomalous ones in the feature representation space (see Figure 1(c)).
+CARLA’s implementation is publicly available on Github: https://github.com/zamanzadeh/CARLA.
   • 2.
@@
 2 Related Work
-In this section, we provide a brief survey of the existing literature that intersects with our research.
-We concentrate on three areas: deep learning for anomaly detection in time series, unsupervised representation of time series, and the contrastive representation learning technique.
+In this section, we concentrate on three areas: deep learning methods in TSAD, unsupervised representation of time series, and the contrastive representation learning technique.
 2.1 Time Series Anomaly Detection
-The detection of anomalies within time series data has been the subject of extensive research, using an array of techniques from statistical methods to classical machine learning and, more recently, deep learning models Schmidl et al.
-(2022).
-Established statistical techniques such as moving averages like the ARIMA model Box and Pierce (1970) have seen widespread application.
-Machine learning techniques, including clustering algorithms and density-based approaches, alongside algorithms similar to decision trees Fei Tony Liu and Zhou (2008) have also been leveraged.
-Deep Learning methods in TSAD: In recent years, deep learning has proven highly effective due to its ability to autonomously extract features Pang et al.
-(2021).
-The focus within TSAD is largely on unsupervised Hundman et al.
-(2018), semi-supervised Park et al.
-(2018), and recently self-supervised Yue et al.
-(2022) approaches, addressing the issue of scarce labelled data.
-Techniques such as OmniAnomaly Su et al.
-(2019) prove to be particularly useful in situations where anomaly labels are unavailable, while semi-supervised methods make efficient use of the labels that are available.
-Deep learning techniques, encompassing autoencoders Sakurada and Yairi (2014); Yao et al.
-(2023), VAEs Park et al.
-(2018), RNNs Su et al.
-(2019), LSTM networks Hundman et al.
-(2018), GANs Li et al.
-(2019), and Transformers Tuli et al.
-(2022); Xu et al.
-(2021a) have shown potential in spotting anomalies within time series, particularly for high-dimensional or non-linear data.
-The preference for deep models like LSTM-VAE Park et al.
-(2018), DITAN Giannoulis et al.
-(2023) and THOC Shen et al.
-(2020) is because they excel at minimising forecasting errors while capturing time series data’s temporal dependencies.
-In summary, semi-supervised methods, such as LSTM-VAE Park et al.
-(2018) excel when labels are readily available.
-On the other hand, unsupervised methods, such as OmniAnomaly Su et al.
-(2019) and AnomalyTransformer Xu et al.
-(2021a), become more suitable when obtaining anomaly labels is a challenge.
+The detection of anomalies within time series data has been the subject of extensive research, using an array of techniques from statistical methods to classical machine learning and, more recently, deep learning models [25].
+Established statistical techniques such as moving averages like the ARIMA model [3] have seen widespread application.
+Machine learning techniques, including clustering algorithms and density-based approaches, alongside algorithms similar to decision trees [8] have also been leveraged.
+Deep Learning methods in TSAD: In recent years, deep learning has proven highly effective due to its ability to autonomously extract features [22].
+The focus within TSAD is largely on unsupervised [12], semi-supervised [23], and recently self-supervised [38] approaches, addressing the issue of scarce labelled data.
+Techniques such as OmniAnomaly [27] prove to be particularly useful in situations where anomaly labels are unavailable, while semi-supervised methods make efficient use of the labels that are available.
+Deep learning techniques, encompassing autoencoders [37], variational autoencoders (VAEs) [23], RNNs [27], LSTM networks [12], GANs [17], and Transformers [29, 34] have shown potential in TSAD, particularly for high-dimensional or non-linear data.
+The preference for deep models like LSTM-VAE [23], DITAN [10] and THOC [26] is because they excel at minimising forecasting errors while capturing time series data’s temporal dependencies.
+In summary, semi-supervised methods, such as LSTM-VAE [23] excel when labels are readily available.
+On the other hand, unsupervised methods, such as OmniAnomaly [27] and AnomalyTransformer [34], become more suitable when obtaining anomaly labels is a challenge.
 These unsupervised deep learning methods are preferred due to their capability to learn robust representations without requiring labelled data.
-The advent of self-supervised learning-based methods has further improved generalization ability in unsupervised anomaly detection Zhao et al.
-(2020).
+The advent of self-supervised learning methods has further improved generalisation in unsupervised anomaly detection [40].
 2.2 Unsupervised Time Series Representation
-The demonstration of substantial performance by unsupervised representation learning across a wide range of fields, such as computer vision Chen et al.
-(2020); Milbich et al.
-(2020), natural language processing Logeswaran and Lee (2018), and speech recognition Xu et al.
-(2021b), has made it a desirable technique.
-In the realm of time series, methods like TKAE Bianchi et al.
-(2019) and TST Zerveas et al.
-(2021) have been suggested.
+The demonstration of substantial performance by unsupervised representation learning across a wide range of fields, such as computer vision [4, 20], natural language processing [18], and speech recognition [35], has made it a desirable technique.
+In the realm of time series, methods like TKAE [2] and TST [39] have been suggested.
 While these methods have made significant contributions, some of them face challenges with scalability for exceptionally long time series or encounter difficulties in modelling complex time series.
-To overcome these limitations, methods such as TNC Tonekaboni et al.
-(2021) and T-Loss Franceschi et al.
-(2019) have been proposed.
-They leverage time-based negative sampling, triplet loss, and local smoothness of signals for the purpose of learning multivariate time series representations that are scalable.
+To overcome these limitations, methods such as TNC [28] and T-Loss [9] have been proposed.
+They leverage time-based negative sampling, triplet loss, and local smoothness of signals for the purpose of learning scalable MTS representations.
 Despite their merits, these methods often limit their universality by learning representations of particular semantic tiers, relying on significant assumptions regarding invariance during transformations.
-Recent studies, such as RoSAS Xu et al.
-(2023), have advanced outlier detection in datasets beyond time series by utilising triplet loss.
+Recent studies, such as RoSAS [33], have outlier detection in datasets beyond time series by utilising triplet loss.
 2.3 Contrastive Representation Learning
-Contrastive representation learning creates an embedding space where similar samples are close and dissimilar ones are distant, applicable in domains like natural language processing, computer vision, and time series anomaly detection Le-Khac et al.
-(2020).
-Traditional models like InfoNCE loss Oord et al.
-(2018), SimCLR Chen et al.
-(2020), and MoCo He et al.
-(2020) use positive-negative pairs to optimize this space, demonstrating strong performance and paving the way for advanced techniques.
-In time series anomaly detection, contrastive learning is crucial for recognising patterns.
-TS2Vec Yue et al.
-(2022) uses this approach hierarchically for multi-level semantic representation, while DCdetector Yang et al.
-(2023) introduces a dual attention asymmetric design with pure contrastive loss, enabling permutation invariant representations.
+Contrastive representation learning creates an embedding space where similar samples are close and dissimilar ones are distant, applicable in domains like natural language processing, computer vision, and time series anomaly detection [16].
+Traditional models like InfoNCE loss [21] and SimCLR [4] use positive-negative pairs to optimise this space, demonstrating strong performance and paving the way for advanced techniques.
+In TSAD, contrastive learning is crucial for recognising patterns.
+TS2Vec [38] uses this approach hierarchically for multi-level semantic representation, while DCdetector [36] introduces a dual attention asymmetric design with pure contrastive loss, enabling permutation invariant representations.
 3 CARLA
@@
 The technique of anomaly injection, a powerful data augmentation strategy for time series, facilitates the application of self-supervised learning in the absence of ground-truth labels.
-This technique has recently been applied in TSAD, notably in COUTA Xu et al.
-(2022).
-While the augmentation methods we employ are not designed to represent every conceivable anomaly type Xu et al.
-(2021a) — a goal that would be unattainable — they amalgamate various robust and generic heuristics to effectively identify prevalent out-of-distribution instances.
+This technique has recently been applied in TSAD, notably in COUTA [32].
+While the augmentation methods we employ are not designed to represent every conceivable anomaly type [34] — a goal that would be unattainable — they amalgamate various robust and generic heuristics to effectively identify prevalent out-of-distribution instances.
 3.1.1 Anomaly Injection Steps
@@
 It is important to underscore that the anomaly injection strategies employed in our model’s evaluation were consistently applied across all benchmarks to ensure a fair and impartial comparison of the model’s performance across various datasets.
-The detailed step-by-step process for our anomaly injection approach is elaborated in Algorithm 1.
-This comprehensive algorithm encapsulates the methodology of both point anomaly injection and subsequence anomaly injection, providing a clear understanding of our process and its implementation.
+The step-by-step process for our anomaly injection approach is detailed in Algorithm 1.
+This comprehensive algorithm encapsulates the methodology of point and subsequence anomaly injection, providing a clear understanding of our process and its implementation.
 Algorithm 1 InjectAnomaly($w$)
@@
 18:       $f\leftarrow\text{random number from }\{\frac{1}{3},\frac{1}{2},2,3\}$triangleright frequency coefficient for Seasonal anomaly
-19:        $w^{\prime}(t)=\begin{cases}w(s+\left(\left\lfloor(t-s)\cdot k\right\rfloor\mod n\right))&\text{if }s\leq t<e\text{ and }k>1\\ w(s+\left\lfloor(t-s)\cdot k\right\rfloor)&\text{if }s\leq t<e\text{ and }0<k<1\\ w(t)&\text{otherwise}\end{cases}$
+19:       $w^{\prime}(t)=\begin{cases}w(s+\left(\left\lfloor(t-s)\cdot f\right\rfloor\mod n\right))&\text{if }s\leq t<e\text{ and }f>1\\ w(s+\left\lfloor(t-s)\cdot f\right\rfloor)&\text{if }s\leq t<e\text{ and }0<f<1\\ w(t)&\text{otherwise}\end{cases}$
 20:    else if $anomaly=\texttt{Trend}$ then
@@
 Our methodology incorporates the injection of single-point anomalies (a spike) at randomly chosen instances within a given window $w_{i}$.
-We employ two distinct types of point anomalies, namely, Global (Figure 2(b)) and Contextual (Figure 2(c)Xu et al.
-(2021a).
+We employ two distinct types of point anomalies, namely, Global (Figure 3(b)) and Contextual (Figure 3(c)[34].
 3.1.3 Subsequence Anomalies
-The technique of subsequence anomaly injection, motivated by the successful implementation of the Outlier Exposure approach Hendrycks et al.
-(2019), enhances anomaly detection in time series data by generating contextual out-of-distribution examples.
+The technique of subsequence anomaly injection, motivated by the successful implementation of the Outlier Exposure approach [11], enhances anomaly detection in time series data by generating contextual out-of-distribution examples.
 It involves the introduction of a subsequence anomaly within a time series window $w_{i}$, represented as $w(t|t\in[s,e])$, where $s$ and $e$ denotes the anomaly’s starting and end time points.
 We explore three primary types of subsequence (a.k.a.
-pattern) anomalies: Seasonal (Figure 2(d)), Shapelet (Figure 2(e)), and Trend (Figure 2(f)Xu et al.
-(2021a).
+pattern) anomalies: Seasonal (Figure 3(d)), Shapelet (Figure 3(e)), and Trend (Figure 3(f)[34].
 (a) Original Window
@@
 Figure 3: Different types of synthetic anomaly injection used in CARLA.
-The figure presents the effect of synthetic anomaly injections into a randomly selected window of size 400 from the first dimension of entity E-2 in the MSL dataset (Hundman et al., 2018).
-The (a) represents the original time series window, while the remaining five demonstrate the same window but with different types of anomalies injected.
-The anomalies are categorised into (b) Global, (c) Contextual, (d) Seasonal, (e) Shapelet, and (f) Trend.
+The figure presents the effect of synthetic anomaly injections into a randomly selected window of size 400 from the first dimension of entity E-2 in the MSL dataset [12].
+The (3(a)) represents the original time series window, while the remaining five demonstrate the same window but with different types of anomalies injected.
+The anomalies are categorised into (3(b)) Global, (3(c)) Contextual, (3(d)) Seasonal, (3(e)) Shapelet, and (3(f)) Trend.
 In each of the five plots with anomalies, the anomalous points or subsequences are accentuated with a red colour.
-The synthetic anomalies were introduced using one or more randomly selected anomaly injection methods, applied sequentially across one or more dimensions of the multivariate time series.
 3.2 Pretext Stage
-The Pretext Stage of CARLA consists of two parts: Part One is a contrastive representation learning step that uses a ResNet architecture (which has shown effectiveness in times series classification tasks Ismail Fawaz et al.
-(2019)) to learn representations for time series windows, and Part Two is a post-processing step that uses the learned representations from Part One to identify semantically meaningful nearest and furthest neighbours for each learned window representation.
+The Pretext Stage of CARLA consists of two parts: Part One is a contrastive representation learning step that uses a ResNet architecture (which has shown effectiveness in times series classification tasks [13]) to learn representations for time series windows, and Part Two is a post-processing step that uses the learned representations from Part One to identify semantically meaningful nearest and furthest neighbours for each learned window representation.
 Algorithm 2 shows the Pretext steps.
@@
 Since the majority of data is normal, the model captures temporal relationships of normal data through learning similar representations for normal windows that are temporally proximate.
-Furthermore, by introducing anomalies into the system, the model can learn a more effective decision boundary, resulting in a reduced false positive rate and enhanced precision in identifying anomalies when compared to current state-of-the-art models.
+Furthermore, by introducing anomalies into the system, the model can learn a more effective decision boundary, resulting in a reduced false positive rate and enhanced precision in anomaly detection compared to current state-of-the-art models.
 The output of Part One includes a trained neural network (ResNet) $\phi_{p}$ and a list of anchor and negative samples stored in $\mathcal{B}$.
 3.2.2 Part Two: Nearest and Furthest Neighbours
-Part Two of our approach uses the feature representations learned in Part One to identify semantically meaningful nearest and furthest neighbours for each sample.
+Part Two of our approach uses the feature representations learned in Part One to identify semantically meaningful nearest and furthest neighbours of each sample.
 To achieve this, we utilise $\mathcal{B}$ along with the indices of $Q$ nearest and $Q$ furthest neighbours of all samples in $\mathcal{B}$.
@@
 $\mathcal{L}_{inconsistency}(\phi_{s},\mathcal{B},\mathcal{F})=-\frac{1}{\lvert\mathcal{B}\rvert}\sum_{w\in\mathcal{B}}\sum_{w_{n}\in\mathcal{F}_{w}}\log(\text{similarity}(\phi_{s},w,w_{n}))$ (4)
-To encourage class diversity and prevent overfitting, we use entropy loss on the distribution of anchor and neighbour samples across classes.
+To encourage class diversity and prevent overfitting, we apply entropy loss on the distribution of anchor and neighbour samples across classes.
 Assuming the classes set is denoted as $\mathcal{C}=\{1,...,C\}$, and the probability of window $w_{i}$ being assigned to class $c$ is denoted as $\phi_{s}^{c}(w_{i})$:
@@
 $\mathcal{L}_{consistency}(\phi_{s},\mathcal{B},\mathcal{N})-\mathcal{L}_{inconsistency}(\phi_{s},\mathcal{B},\mathcal{F})-\beta\cdot\mathcal{L}_{entropy}(\phi_{s},\mathcal{B},\mathcal{C})$
-The goal of the loss function is to learn a representation that is highly discriminative, with nearest neighbours assigned to the same class as the anchor samples, and furthest neighbours assigned to a distinct class.
+The goal of the loss function is to learn a representation that is highly discriminative, with the nearest neighbours assigned to the same class as the anchor samples and the furthest neighbours assigned to a distinct class.
 By incorporating the semantically meaningful nearest and furthest neighbours, the model is able to produce more consistent and confident predictions, with the probability of a window being classified as one particular class is close to 1 or 0.
@@
 Table 1: Statistics of the benchmark datasets used.
-Benchmark # datasets # dims Train size Test size Anomaly ratio
+Benchmark # datasets # dims Train size Test size Anomaly%
 MSL 27 55 58,317 73,729 10.72%
@@
 WADI 1 123 784,568 172,801 5.77%
-Yahoo - A1 Benchmark 67 1 46,667 46,717 1.76%
+Yahoo-A1 67 1 46,667 46,717 1.76%
 KPI 29 1 1,048,576 2,918,847 1.87%
@@
 The datasets are summarised in Table 1, which provides key statistics for each dataset.
-NASA DatasetsMars Science Laboratory (MSL) and Soil Moisture Active Passive (SMAP) [^1] https://www.kaggle.com/datasets/patrickfleith/nasa-anomaly-detection-dataset-smap-msl – (Hundman et al., 2018) are real-world datasets collected from NASA spacecraft.
-These datasets contain anomaly information derived from reports of incident anomalies for a spacecraft monitoring system.
-MSL and SMAP comprises of 27 and 55 datasets respectively and each equipped with a predefined train/test split, where, unlike other datasets, their training set are unlabeled.
-Server Machine Dataset (SMD) [^2] https://github.com/NetManAIOps/OmniAnomaly/tree/master/ServerMachineDataset (Su et al., 2019) is gathered from 28 servers, incorporating 38 sensors, over a span of 10 days.
+All datasets, except Yahoo, have a predefined train/test split with unlabeled training data.
+NASA DatasetsMars Science Laboratory (MSL) and Soil Moisture Active Passive (SMAP) [^1] https://www.kaggle.com/datasets/patrickfleith/nasa-anomaly-detection-dataset-smap-msl – [12] These real-world datasets, collected from NASA spacecraft, contain anomaly information from incident reports for a spacecraft monitoring system.
+Server Machine Dataset (SMD) [^2] https://github.com/NetManAIOps/OmniAnomaly/tree/master/ServerMachineDataset [27] is gathered from 28 servers, incorporating 38 sensors, over a span of 10 days.
 During this period, normal data was observed within the initial 5 days, while anomalies were sporadically injected during the subsequent 5 days.
-The dataset is also equipped with a predefined train/test split, where the training data is unlabeled.
-Secure water treatment (SWaT) [^3] https://itrust.sutd.edu.sg/testbeds/secure-water-treatment-swat/ (Mathur and Tippenhauer, 2016) is meticulously gathered within 11 days, utilising a water treatment platform that consisted of 51 sensors.
-During the initial 7 days, the dataset exclusively comprised normal data, accurately reflecting regular operational conditions.
-However, in the subsequent 4 days, a total of 41 anomalies were deliberately generated using a wide range of attack methods.
-Water distribution testbed (WADI) [^4] https://itrust.sutd.edu.sg/testbeds/water-distribution-wadi/ (Ahmed et al., 2017) is acquired from a scaled-down urban water distribution system that included a combined total of 123 actuators and sensors.
+Secure water treatment (SWaT) [^3] https://itrust.sutd.edu.sg/testbeds/secure-water-treatment-swat/ [19] is meticulously gathered over 11 days, utilising a water treatment platform with 51 sensors.
+During the initial 7 days, it consisted exclusively of normal data, reflecting regular operational conditions.
+In the following 4 days, 41 anomalies were deliberately generated using a wide range of attacks.
+Water distribution testbed (WADI) [^4] https://itrust.sutd.edu.sg/testbeds/water-distribution-wadi/ [1] is acquired from a scaled-down urban water distribution system that included a combined total of 123 actuators and sensors.
 This data covers a span of 16 days.
@@
 In addition, the test dataset encompasses 15 distinct anomaly segments, providing valuable examples for anomaly detection and analysis purposes.
-Yahoo [^5] https://webscope.sandbox.yahoo.com/catalog.php?datatype=s&did=70 dataset (Laptev et al., 2015) is a benchmark dataset that consists of hourly 367 sampled time series with anomaly points labels.
+Yahoo [^5] https://webscope.sandbox.yahoo.com/catalog.php?datatype=s&did=70 dataset [15] is a benchmark dataset that consists of hourly 367 sampled time series with anomaly points labels.
 The dataset includes four benchmarks, with the A1 benchmark containing ”real” production traffic data from Yahoo properties, and the other three are synthetic.
-For evaluation purposes, we focus on the A1 benchmark, which includes 67 univariate time series.
-However, some of these contain no anomalies in the test, hence we exclude them from the evaluation.
-Key performance indicators (KPI) [^6] https://github.com/NetManAIOps/KPI-Anomaly-Detection contains time series data of service and machine key performance indicators (KPIs) from real scenarios of Internet companies.
-Service KPIs measure web service performance, while machine KPIs reflect machine health.
-Examples of service KPIs include page response time and page views, while machine KPIs include CPU and memory utilisation.
-The dataset includes 29 different KPI datasets, with separate training and test datasets provided.
+For evaluation, we focus on the A1 benchmark, which includes 67 univariate time series.
+We use half of each dataset for training and the remaining for test, excluding those without anomalies in the test set.
+Key performance indicators (KPI) [^6] https://github.com/NetManAIOps/KPI-Anomaly-Detection contains data of service and machine key performance indicators (KPIs) from real scenarios of Internet companies.
+Service KPIs measure web service performance, such as page response time and page views.
+Machine KPIs reflect machine health, including CPU and memory utilisation.
 4.2 Benchmark Methods
-In this overview, provide a description of the ten prominent TSAD models that were used for comparison with CARLA, spanning various TSAD categories.
-For the semi-supervised approach, the LSTM-VAE (Park et al., 2018) is highlighted, which is a point-wise anomaly detection model.
-In the realm of unsupervised reconstruction-based models, Donut (Xu et al., 2018) stands out for UTS, while OmniAnomaly Su et al.
-(2019) and AnomalyTransformer (Xu et al., 2021a) are noted for MTS, alongside TranAD (Tuli et al., 2022) which caters to both UTS and MTS.
-The unsupervised forecasting-based model, THOC Shen et al.
-(2020), is recognized for its predictive capabilities.
-Within the self-supervised category, MTAD-GAT (Zhao et al., 2020) emerges as a hybrid model, with TS2Vec Yue et al.
-(2022) and DCdetector Yang et al.
-(2023) offering representation-based methodologies.
-Additionally, recent advancements are represented by TimesNet Wu et al.
-(2023) and DCdetector Yang et al.
-(2023), showcasing the ongoing evolution in TSAD models.
-Donut  [^7] https://github.com/NetManAIOps/donut Xu et al.
-(2018) is a method for detection anomalies in UTS.
-It utilises a VAE for a reconstruction-based strategy.
-By learning the latent representation of time series data, it accurately identifies anomalies when data reconstruction significantly deviates from the original.
-LSTM-VAE [^8] https://github.com/TimyadNyda/Variational-Lstm-Autoencoder Park et al.
-(2018) is a point-wise model that combines LSTM networks, which capture temporal data dependencies, with VAEs for generating high-dimensional data distributions and modelling complex data patterns.
-This integration allows for effective anomaly detection in MTS points by identifying deviations from normal behaviour, which is crucial in applications like robot-assisted feeding for ensuring safety and efficiency.
-OmniAnomaly [^9] https://github.com/smallcowbaby/OmniAnomaly Su et al.
-(2019) is an unsupervised model that utilises a Variational Autoencoder (VAE) to model multivariate time series data.
-By analyzing the likelihood of a given window’s reconstruction, the model can detect anomalies.
-MTAD-GAT [^10] https://github.com/mangushev/mtad-gat Zhao et al.
-(2020) enhances MTS anomaly detection by leveraging graph attention networks to understand complex dependencies across different time series.
-It employs a self-supervised framework that uses graph attention layers to learn relationships in both temporal and feature dimensions.
-This approach optimizes both forecasting-based and reconstruction-based models for anomaly detection.
-THOC [^11] We utilised the authors’ shared implementation, as it is not publicly available.
-Shen et al.
-(2020) utilises a multi-layer dilated recurrent neural network (RNN) alongside skip connections in order to effectively handle contextual information.
-THOC utilise a temporal hierarchical one-class network for anomaly detection.
-AnomalyTransformer [^12] https://github.com/thuml/Anomaly-Transformer Xu et al.
-(2021a) introduces an approach for detecting anomalies in MTS by leveraging the Transformer architecture, known for its effectiveness in capturing long-range dependencies.
-It employs association discrepancy to identify anomalies, focusing on the differences in relationships among data points in a time series.
-TranAD [^13] https://github.com/imperial-qore/TranAD Tuli et al.
-(2022) leverages deep transformer networks for efficient anomaly detection in MTS and UTS.
-It employs attention-based sequence encoders for quick inference, leveraging focus score-based self-conditioning and adversarial training for robust feature extraction and stability.
-TS2Vec [^14] https://github.com/yuezhihan/ts2vec Yue et al.
-(2022) is an unsupervised contrastive model that is capable of learning multiple contextual representations of MTS and UTS semantically at various levels.
-TS2Vec employs contrastive learning in a hierarchical way, which provides a contextual representation.
-They proposed a method to be employed for time series anomaly detection.
-DCdetector [^15] https://github.com/DAMO-DI-ML/KDD2023-DCdetector Yang et al.
-(2023) introduces an approach for time series anomaly detection by employing dual attention mechanisms and contrastive representation learning.
-This method focuses on capturing intricate temporal relationships and distinguishing between normal and anomalous patterns in time series data.
-TimesNet [^16] https://github.com/thuml/TimesNet Wu et al.
-(2023) proposes a framework for analyzing general time series data by modelling temporal 2D variations.
-This approach allows for capturing complex patterns and dependencies in time series.
-This method leverages advanced neural networks to enhance analysis accuracy and efficiency across various applications like time series anomaly detection.
-Random Anomaly Score.
-The “Random Anomaly Score” model is designed to generate anomaly scores for time series windows based on a normal distribution ($\mathcal{N}(0,1)$) with mean 0 and standard deviation 1.
-This model has been specifically developed to showcase evaluation metrics and strategies in the field of TSAD.
+We provide a description of the ten prominent TSAD models that were used for comparison with CARLA, spanning various TSAD categories.
+For the semi-supervised approach, the LSTM-VAE [23] is highlighted, which is a point-wise anomaly detection model.
+In the realm of unsupervised reconstruction-based models, Donut [31] stands out for UTS, while OmniAnomaly [27] and AnomalyTransformer [34] are noted for MTS, alongside TranAD [29] which caters to both UTS and MTS.
+The unsupervised forecasting-based model, THOC [26] and TimesNet [30], are recognised for their predictive capabilities.
+Within the self-supervised category, MTAD-GAT [40] emerges as a hybrid model, with TS2Vec [38] and DCdetector [36] offering representation-based methodologies.
+Additionally, recent advancements are represented by TimesNet [30] and DCdetector [36], showcasing the ongoing evolution in TSAD models.
+The “Random Anomaly Score” model is designed to generate anomaly scores based on a normal distribution ($\mathcal{N}(0,1)$) with mean 0 and standard deviation 1.
+This model is specifically developed to illustrate evaluation metrics and strategies in the field of TSAD.
 4.3 Evaluation Setup
-In our study, we evaluate ten TSAD models on benchmark datasets previously mentioned in section 4.1 using their best hyper-parameters as they stated to ensure a fair evaluation.
+In our study, we evaluate ten TSAD models on benchmark datasets previously mentioned in section 4.1 using their best hyper-parameters, as they stated, to ensure a fair evaluation.
 We summarise the default hyper-parameters used in our implementation as follows: The CARLA model consists of a 3-layer ResNet architecture with three different kernel sizes [8, 5, 3] to capture temporal dependencies.
@@
 It is important to note that we do not use Point Adjustment (PA) in our evaluation process.
-Despite its popularity, we do not use PA due to the findings in Kim et al.
-(2022) that its application leads to an overestimation of a TSAD model’s capability and can bias results in favour of methods that produce extreme anomaly scores.
-Thus, to ensure the accuracy and validity of our results, we relegate PA results to B and present here conventional F1 scores.
-The F1 score calculated without considering the PA is referred to as F1, while the adjusted predictions are denoted as $F1_{PA}$ in B.
+Despite its popularity, [14] found that applying PA leads to an overestimation of TSAD models’ capability and can bias results towards methods that produce extreme anomaly scores.
+To ensure accuracy, we present conventional F1 scores and relegate PA results to B.
+The F1 score without PA is referred to as F1.
 Table 2: Precision (Prec), recall (Rec), F1 and AU-PR results for various models on multivariate time series datasets.
-The most optimal evaluation results are displayed in bold, while the second best ones are indicated by underline.
+The best results are in bold, and the second-best is indicated by underline.
 Due to the single time series in both the SWaT and WADI datasets, the standard deviation of AU-PR is not available.
 Model Metric MSL SMAP SMD SWaT WADI Avg Rank
-LSTM-VAE Park et al. (2018) (2018) Prec 0.2723 0.2965 0.2045 0.9707 0.0596 5.5
+LSTM-VAE [23] (2018) Prec 0.2723 0.2965 0.2045 0.9707 0.0596 5.5
 Rec 0.8083 0.8303 0.5491 0.5769 1.0000
@@
 AU-PR 0.285$\pm$0.249 0.258$\pm$0.305 0.395$\pm$0.257 0.685 0.039
-OmniAnomaly Su et al. (2019) (2019) Prec 0.1404 0.1967 0.3067 0.9068 0.1315 5.0
+OmniAnomaly [27] (2019) Prec 0.1404 0.1967 0.3067 0.9068 0.1315 5.0
 Rec 0.9085 0.9424 0.9126 0.6582 0.8675
@@
 AU-PR 0.149$\pm$0.182 0.115$\pm$0.129 0.365$\pm$0.202 0.713 0.120
-MTAD-GAT Zhao et al. (2020) (2020) Prec 0.3559 0.3783 0.2473 0.1387 0.0706 4.2
+MTAD-GAT [40] (2020) Prec 0.3559 0.3783 0.2473 0.1387 0.0706 4.2
 Rec 0.7067 0.8239 0.5834 0.9585 0.5838
@@
 AU-PR 0.335$\pm$0.259 0.339$\pm$0.300 0.401$\pm$0.263 0.095 0.084
-THOC Shen et al. (2020) (2020) Prec 0.1936 0.2039 0.0997 0.5453 0.1017 6.8
+THOC [26] (2020) Prec 0.1936 0.2039 0.0997 0.5453 0.1017 6.8
 Rec 0.7718 0.8294 0.5307 0.7688 0.3507
@@
 AU-PR 0.239$\pm$0.273 0.195$\pm$0.262 0.107$\pm$0.126 0.537 0.103
-AnomalyTransformer Xu et al. (2021a) (2021) Prec 0.2182 0.2669 0.2060 0.9711 0.0601 5.4
+AnomalyTran [34] (2021) Prec 0.2182 0.2669 0.2060 0.9711 0.0601 5.4
 Rec 0.8231 0.8600 0.5822 0.5946 0.9604
@@
 AU-PR 0.236$\pm$0.237 0.264$\pm$0.315 0.273$\pm$0.232 0.681 0.040
-TranAD Tuli et al. (2022) (2022) Prec 0.2957 0.3365 0.2649 0.1927 0.0597 6.2
+TranAD [29] (2022) Prec 0.2957 0.3365 0.2649 0.1927 0.0597 6.2
 Rec 0.7763 0.7881 0.5661 0.7965 1.0000
@@
 AU-PR 0.278$\pm$0.239 0.287$\pm$0.300 0.412$\pm$0.260 0.192 0.039
-TS2Vec Yue et al. (2022) (2022) Prec 0.1832 0.2350 0.1033 0.1535 0.0653 7.7
+TS2Vec [38] (2022) Prec 0.1832 0.2350 0.1033 0.1535 0.0653 7.7
 Rec 0.8176 0.8826 0.5295 0.8742 0.7126
@@
 AU-PR 0.132$\pm$0.135 0.148$\pm$0.165 0.113$\pm$0.075 0.136 0.057
-DCdetector Yang et al. (2023) (2023) Prec 0.1288 0.1606 0.0432 0.1214 0.1417 9.1
+DCdetector [36] (2023) Prec 0.1288 0.1606 0.0432 0.1214 0.1417 9.1
 Rec 0.9578 0.9619 0.9967 0.9999 0.9684
@@
 AU-PR 0.129$\pm$0.144 0.124$\pm$0.153 0.043$\pm$0.036 0.126 0.121
-TimesNet Wu et al. (2023) (2023) Prec 0.2257 0.2587 0.2450 0.1214 0.1334 6.1
+TimesNet [30] (2023) Prec 0.2257 0.2587 0.2450 0.1214 0.1334 6.1
 Rec 0.8623 0.8994 0.5474 1.0000 0.1565
@@
 AU-PR 0.501$\pm$0.267 0.448$\pm$0.326 0.507$\pm$0.195 0.681 0.126
-Table 3: Precision (Prec), recall (Rec), F1, and AU-PR results for various models on univariate time series datasets.
-The most optimal evaluation results are displayed in bold, while the second best ones are indicated by underline.
+Table 3: Precision (Prec), recall (Rec), F1, and AU-PR results for various models on univariate time series datasets. The best results are in bold, and the second-best is indicated by underline.
 Model Metric Yahoo-A1 KPI Avg Rank
-Donut Xu et al. (2018) (2018) Prec 0.3239 0.0675 5.0
+Donut [31] (2018) Prec 0.3239 0.0675 5.0
 Rec 0.9955 0.9340
@@
 AU-PR 0.264$\pm$0.244 0.078$\pm$0.073
-THOC Shen et al. (2020) (2020) Prec 0.1495 0.1511 5.5
+THOC [26] (2020) Prec 0.1495 0.1511 5.5
 Rec 0.8326 0.5116
@@
 AU-PR 0.349$\pm$0.342 0.229$\pm$0.217
-TranAD Tuli et al. (2022) (2022) Prec 0.4185 0.2235 2.0
+TranAD [29] (2022) Prec 0.4185 0.2235 2.0
 Rec 0.8712 0.4016
@@
 AU-PR 0.691$\pm$0.324 0.285$\pm$0.206
-TS2Vec Yue et al. (2022) (2022) Prec 0.3929 0.1333 5.0
+TS2Vec [38] (2022) Prec 0.3929 0.1333 5.0
 Rec 0.6305 0.4329
@@
 AU-PR 0.491$\pm$0.352 0.221$\pm$0.156
-DCdetector Yang et al. (2023) (2023) Prec 0.0598 0.0218 8.0
+DCdetector [36] (2023) Prec 0.0598 0.0218 8.0
 Rec 0.9434 0.8589
@@
 AU-PR 0.041$\pm$0.059 0.018$\pm$0.017
-TimesNet Wu et al. (2023) (2023) Prec 0.3808 0.2174 3.0
+TimesNet [30] (2023) Prec 0.3808 0.2174 3.0
 Rec 0.7883 0.2713
@@
 Additionally, we computed the average ranks of the models based on their F1.
-For all methods, we used the precision-recall curve on the anomaly score for each time series in the benchmark datasets to find the best F1 score based on precision and recall for the target time series.
+For all methods, we used the precision-recall curve on the anomaly score for each time series in the datasets to find the best F1 score based on precision and recall for the target time series.
 Since certain benchmark datasets such as MSL contain multiple time series datasets (as shown in Table 1), we cannot merge or combine these time series due to the absence of timestamp information.
@@
 Furthermore, for datasets with more than one time series, we report the average and standard deviation of AU-PR across all time series.
-AU-PR is advantageous in imbalanced datasets because it is less sensitive to the distribution of classes Saito and Rehmsmeier (2015).
-It considers the trade-off between precision and recall across all possible decision thresholds, making it more robust in scenarios where the number of instances in classes are imbalanced.
+AU-PR is advantageous in imbalanced datasets because it is less sensitive to the distribution of classes [24].
+It considers the trade-off between precision and recall across all possible decision thresholds, making it more robust in scenarios where the number of instances in classes is imbalanced.
 By focusing on the positive class (i.e.
 anomalous) and its predictions, AU-PR offers a more precise evaluation of a model’s ability to identify and prioritise the minority class correctly.
+Figure 4: Critical difference diagrams for multivariate (4(a)) and univariate (4(b)) time series benchmark datasets.
 Tables 2 and 3 demonstrate the performance comparison between CARLA and all benchmark methods for multivariate and univariate datasets, respectively.
@@
 While the recall is higher in other benchmarks, such as OmniAnomaly, it is with the expense of very low precision (with median precision $<19.67\%$ across all datasets).
-This means that a good balance is achieved between precision and recall by CARLA.
+This means that a good balance between precision and recall is achieved by CARLA.
 This is also shown in our consistently high AU-PR compared to other models and indicates that it is proficient in accurately identifying anomalous instances with higher recall while minimising false alarms with higher precision on imbalanced datasets where normal instances are predominant.
 Finally, CARLA achieved the lowest average rank (best rank), based on F1, in both multivariate and univariate benchmarks (see last column in Tables 2 and 3).
-Figure 4: Models’ performance comparison on multivariate time series datasets regarding their FPR and AU-PR.
-Models in the optimal quadrant (low FPR, high AU-PR) demonstrate superior anomaly detection with minimal false alarms.
-Figure 5: Models’ performance comparison on univariate time series datasets regarding their FPR and AU-PR.
-Models in the optimal quadrant (low FPR, high AU-PR) demonstrate superior anomaly detection with minimal false alarms.
+Figure 4 shows the average ranking for each model in a critical difference diagram [6].
+Models within the same clique (black bar) are not statistically significant.
+Statistically, the Wilcoxon signed-rank test with Holm correction is used as a post hoc test to Friedman’s test.
+The diagrams cover 208 univariate and multivariate datasets (detailed in Table 1).
+CARLA consistently ranks as the best-performing model, scoring 1.00 in univariate datasets (Figure 4(b)) and 1.60 in multivariate datasets (Figure 4(a)), highlighting its superiority.
+Moreover, CARLA is significantly different from other models.
+Figure 5: Models’ performance comparison on MTS & UTS datasets regarding their FPR and AU-PR.
+Models in the optimal quadrant (low FPR, high AU-PR) have superior anomaly detection with minimal false alarms.
 We now evaluate the performance of all models in terms of FPR.
 We also incorporate AU-PR in this comparison, as some models can achieve a low FPR at the expense of lowering AU-PR.
-In the scatter plots provided in Figure 4 we show FPR on the x-axis and AU-PR on the y-axis.
+In the scatter plots provided in Figure 5, we show FPR on the x-axis and AU-PR on the y-axis.
 The ideal point on this scatter plot is in the leftmost top corner.
@@
 This is particularly noteworthy in the context of anomaly detection, where the cost of false positives can be substantial, leading to wasted resources and potential overlook of true anomalies.
-CARLA’s position in the plots are close to leftmost top corner (in the optimal quadrant).
-This indicates fewer false alarms while maintaining competitive AU-PR scores, which measures the precision and recall balance of anomaly detection.
+CARLA’s position in the plots is close to the leftmost top corner (in the optimal quadrant).
+This indicates fewer false alarms while maintaining competitive AU-PR scores, which measure the precision and recall balance of anomaly detection.
 For instance, on the MSL dataset, CARLA is among the models with the lowest FPR, and its AU-PR score is in the upper tier, showcasing its ability to identify true anomalies accurately.
@@
 In the SWaT dataset, CARLA excels with the lowest FPR and the third highest AU-PR, marking it suitable for precision-critical applications.
-Even on WADI, while CARLA’s FPR is slightly higher, it remains in the optimal quadrant, reflecting a strong FPR and AU-PR balance vital for maintaining operational continuity and effective anomaly detection.
+Even on WADI, while CARLA’s FPR is slightly higher, it remains in the optimal quadrant, reflecting a strong FPR and AU-PR balance vital for operational continuity and effective anomaly detection.
 Adding to CARLA’s impressive performance on multivariate datasets, its capabilities in univariate time series analysis, as shown in Figure 5, further highlighting its robustness.
@@
 Its place in the upper left quadrant indicates a strong ability to correctly identify anomalies with a minimal rate of false alarms.
-On the KPI dataset, CARLA maintains a competitive edge, presenting a low FPR that surpasses most other models.
+On the KPI dataset, CARLA maintains a competitive edge with a low FPR that surpasses most other models.
 While the AU-PR is the highest, CARLA’s ability to limit false positives is a notable strength.
@@
 4.5 Ablation Study
-Additionally, for a deeper understanding of the individual contributions of the stages and different components in our proposed model, we conduct an ablation study.
-Our analysis focused on the following aspects: (i) Effectiveness of CARLA’s two stages, (ii) Positive pair selection strategy (iii) Effectiveness of Different Anomaly Types (iv) Effectiveness of the loss components.
+For a deeper understanding of the contributions of different stages and components in our proposed model, we conduct an ablation study.
+Our analysis focused on: (i) Effectiveness of CARLA’s two stages, (ii) Positive pair selection strategy (iii) Effectiveness of Different Anomaly Types (iv) Effectiveness of the loss components.
 4.5.1 Effectiveness of CARLA’s Stages
@@
 For this purpose, we employ t-distributed Stochastic Neighbor Embedding (t-SNE) to visualise the output of the Pretext Stage and the Self-supervised Classification stage.
-These are graphically represented in Figures 5(a) and 5(b) correspondingly.
+These are graphically represented in Figures 6(a) and 6(b) correspondingly.
 From these graphical representations, it is palpable that the second stage significantly enhances the discrimination capacity between normal and anomalous samples.
@@
 Figure 6: Comparative analysis of the model’s stages utilising t-SNE and anomaly score distributions on M-6 dataset in MSL.
-(a) t-SNE after the Pretext stage, indicating the initial representation of normal and anomalous samples.
-(b) t-SNE after the Self-supervised Classification stage, demonstrating enhanced segregation of normal and anomalous instances.
-(c) Anomaly score distribution in the Pretext Stage based on Euclidean distances between test and training samples.
-(d) The distribution of anomaly scores following the application of the Self-supervised Classification Stage, evidencing an improved representation.
+(6(a)) t-SNE after the Pretext stage, indicating the initial representation of normal and anomalous samples.
+(6(b)) t-SNE after the Self-supervised Classification stage, demonstrating enhanced segregation of normal and anomalous instances.
+(6(c)) Anomaly score distribution in the Pretext Stage based on Euclidean distances between test and training samples.
+(6(d)) The distribution of anomaly scores following the application of the Self-supervised Classification Stage evidencing an improved representation.
 Further, we assessed the efficacy of the Self-supervised Classification Stage in refining the representation generated by the Pretext Stage.
@@
 Conversely, Figure 6(d) depicts the subsequent alteration in the distribution of anomaly scores following the application of the Self-supervised Classification Stage.
-In the Pretext Stage, the anomaly score is computed as the minimum distance between the test sample and all original training samples in the representation space.
-This approach aims to identify the closest match among the training samples for the given test sample.
+In the Pretext Stage, the anomaly score is computed as the minimum distance between the test sample and all original training samples in the representation space, aiming to identify the closest match among the training samples for the given test sample.
 A smaller distance implies a greater probability that the test sample belongs to the normal data distribution.
 In the Self-supervised Classification Stage, we use the inference step in Section 3.3 of the article. Equation 8 provides further details on the computation of the anomaly label.
-As we can observe from the figures 5(a) and 5(b), the self-supervised classification Stage has resulted in a significant improvement in the separation of normal and anomalous windows.
+As we can observe from the figures 6(a) and 6(b), the self-supervised classification Stage has resulted in a significant improvement in the separation of normal and anomalous windows.
 The distribution of anomaly scores in Figure 6(d) is more clearly separated than in Figure 6(c).
 This indicates that the second stage has produced more consistent representations.
-Furthermore, the anomaly scores in Figure 5(d) are relatively closer to 0 or 1, which indicates an improvement in discrimination between normal and anomalous windows in the second stage of CARLA.
+Furthermore, the anomaly scores in Figure 6(d) are relatively closer to 0 or 1, indicating improved discrimination between normal and anomalous windows in the second stage of CARLA.
 4.5.2 Positive Pair Selection
@@
 Our experiment demonstrated that selecting positive pairs using a random temporal neighbour is a more effective approach than weak augmentation with noise.
-Since anomalies occur rarely, choosing a positive pair that is in the target window’s temporal proximity is a better strategy in our model.
+Since anomalies occur rarely, choosing a positive pair that is in the target window’s temporal proximity is a better strategy.
 4.5.3 Effectiveness of Different Anomaly Types
-Based on the results of the ablation study on the MSL dataset, CARLA’s performance is evaluated by systematically removing different types of anomalies during the Pretext Stage.
+CARLA’s performance is evaluated on the MSL dataset by systematically removing different types of anomalies during the Pretext Stage.
 The evaluation metrics used are precision, recall, F1 score, and AU-PR.
@@
 However, the recall improves, suggesting that the model becomes more sensitive in detecting anomalies overall.
-Removing seasonal anomalies has a noticeable effect on the model’s precision, which drops significantly.
-The consistently high recall suggests the model effectively captures many anomalies without relying on seasonal patterns.
+Removing seasonal anomalies significantly drops precision, though the high recall indicates effective anomaly detection without relying on seasonal patterns.
 Overall, the analysis of the results suggests that global and seasonal anomalies play a relatively more significant role in the CARLA model’s performance, while contextual, shapelet, and trend anomalies have varying degrees of impact.
@@
 4.6 Parameter Sensitivity
-We also examine the sensitivity of CARLA’s parameters. We investigate and analyze the effect of the window size, number of classes, number of NN/FN, and coefficient of entropy.
+We examine the sensitivity of CARLA’s parameters, investigating and analysing the effect of the window size, number of classes, number of NN/FN, and entropy coefficient.
 4.6.1 Effect of Window Size
@@
 Table 7 and Figure 7 present the results of exploring the impact of window size on three datasets: MSL, SMD, and Yahoo.
-Figure 7: (a) F1 Score and (b) AU-PR for different Window sizes across MSL, SMD, and Yahoo datasets
 Table 7: Exploring the effect of window size on MSL, SMD, and Yahoo datasets.
 MSL SMD Yahoo
-WS
-Prec
-Rec
-F1
-AUPR
-Prec
-Rec
-F1
-AUPR
-Prec
-Rec
-F1
-AUPR
-50
-0.2755
-0.7702
-0.4058
-0.3173
-0.4569
-0.4885
-0.4722
-0.4636
-0.3046
-0.8800
-0.4526
-0.5533
-100
-0.3233
-0.7666
-0.4548
-0.4155
-0.2584
-0.5938
-0.3598
-0.4136
-0.3709
-0.9713
-0.5368
-0.5331
-150
-0.3606
-0.7823
-0.4936
-0.4595
-0.3599
-0.6592
-0.4655
-0.4661
-0.5720
-0.9578
-0.7162
-0.6685
-200
-0.3891
-0.7959
-0.5227
-0.5009
-0.4276
-0.6362
-0.5114
-0.5070
-0.5747
-0.9755
-0.7233
-0.6450
-250
-0.4380
-0.8037
-0.5412
-0.5188
-0.4224
-0.6596
-0.5150
-0.4850
-0.7129
-0.9781
-0.8247
-0.7164
+WS Prec Rec F1 AUPR Prec Rec F1 AUPR Prec Rec F1 AUPR
+50 0.2755 0.7702 0.4058 0.3173 0.4569 0.4885 0.4722 0.4636 0.3046 0.8800 0.4526 0.5533
+100 0.3233 0.7666 0.4548 0.4155 0.2584 0.5938 0.3598 0.4136 0.3709 0.9713 0.5368 0.5331
+150 0.3606 0.7823 0.4936 0.4595 0.3599 0.6592 0.4655 0.4661 0.5720 0.9578 0.7162 0.6685
+200 0.3891 0.7959 0.5227 0.5009 0.4276 0.6362 0.5114 0.5070 0.5747 0.9755 0.7233 0.6450
+250 0.4380 0.8037 0.5412 0.5188 0.4224 0.6596 0.5150 0.4850 0.7129 0.9781 0.8247 0.7164
+Figure 7: (a) F1 Score and (b) AU-PR for different window sizes across MSL, SMD, and Yahoo datasets
 Based on the analysis of both the F1 score and AU-PR, it can be concluded that window size 200 consistently outperforms other window sizes on the MSL, SMD, and Yahoo datasets overall.
 This window size strikes a balance between precision and recall, effectively capturing anomalies while maintaining a high discrimination ability.
-Therefore, window size 200 is selected as the best choice for conducting time series analysis in all experiments.
+Therefore, window size 200 is selected as the best choice in all experiments.
 Additionally, Figure 7 shows that our model’s performance is stable in window sizes between 150 and 250.
-Figure 8: F1 Score and AU-PR for (a) different number of classes, (b) different numbers of neighbours in $\mathcal{N}$$\mathcal{F}$ and (c) different coefficient of entropy, on MSL dataset.
 4.6.2 Effect of Number of Classes
@@
 Based on the illustration of both the F1 score and AU-PR in Figure 8(a), the model performs relatively well across different numbers of classes.
-While CARLA’s performance is pretty stable across all number of classes denoted on the x-axis of the plot in Figure 7(a), it can be concluded that using 10 classes as the number of class parameters yields the highest performance.
-It strikes a balance between capturing anomalies and minimising false positives while effectively discriminating between anomalies and normal data points.
+While CARLA’s performance is pretty stable across all number of classes denoted on the x-axis of the plot in Figure 8(a), it can be concluded that using 10 classes yields the highest performance.
+It strikes a balance between capturing anomalies and minimising false positives while effectively discriminating between anomalies and normal samples.
 The results suggest that increasing the number of classes beyond 10 does not significantly improve the model’s performance.
 When the number of classes increases, normal representations divide into the different classes and, in CARLA, are detected as anomalies (lower probability of belonging to the major class).
-However, using 2 classes leads to a lower performance compared to the selected parameter of 10 classes, implying that a more fine-grained classification with 10 classes provides the assumption that various anomalies are spread in the representation space.
+However, using 2 classes leads to a lower performance compared to 10 classes, implying that a more fine-grained classification with 10 classes provides the assumption that various anomalies are spread in the representation space.
+Figure 8: F1 Score and AU-PR for (a) different number of classes, (b) different numbers of neighbours in $\mathcal{N}$$\mathcal{F}$ and (c) different coefficient of entropy, on MSL dataset.
 4.6.3 Effect of Number of neighbours in $\mathcal{N}$ and $\mathcal{F}$
@@
 This number of parameters strikes a favourable balance between precision and recall, allowing for effective anomaly detection while maintaining a high discrimination ability.
-Therefore, selecting 5 for the number of neighbours is deemed the optimal choice for this experiment, as it consistently achieves higher F1 scores and AU-PR values.
-4.6.4 Effect of Coefficient of Entropy
+Therefore, selecting 5 neighbours is deemed the optimal choice for this experiment, as it consistently achieves higher F1 scores and AU-PR values.
+4.6.4 Effect of Entropy Coefficient
 In this experiment, we explore the impact of the entropy coefficient.
@@
 The coefficient of 0 significantly reduces the model’s performance, further emphasising the importance of incorporating the entropy component for improved anomaly detection performance.
-Overall, the analysis indicates that coefficients of 2, 5, and 10 are viable for the entropy component in the loss function, but a coefficient of 5 is chosen as optimal due to its superior F1 score and AU-PR.
+Overall, the analysis indicates that coefficients of 2, 5, and 10 are viable for the entropy component in the loss function, with 5 being optimal due to its superior F1 score and AU-PR.
 5 Conclusion
-In conclusion, our innovative end-to-end self-supervised framework, CARLA, utilises contrastive representation learning with anomaly injection to generate robust representations for time series windows and classify anomalous windows.
-The use of semantically meaningful nearest and furthest neighbours as a prior allows us to capture underlying patterns in time series data and learn a representation that is well-aligned with these patterns, thereby enhancing the accuracy of anomaly detection.
-Additionally, the experimental evaluation shows promising results in detecting anomalies in the seven most commonly used real-world time series datasets, demonstrating the effectiveness of CARLA in this domain.
-This research has demonstrated the potential of contrastive learning combined with synthetic anomaly injection to overcome the limitations of the lack of labelled data in time series anomaly detection.
+Our innovative end-to-end self-supervised framework, CARLA, utilises contrastive representation learning with anomaly injection to generate robust representations for time series and classify anomalous windows.
+The use of semantically meaningful nearest and furthest neighbours as a prior allows us to capture underlying patterns in time series and learn a representation that is well-aligned with these patterns, thereby enhancing detection accuracy.
+Our extensive experimental evaluation of the seven most commonly used time series benchmarks, encompassing 208 datasets with diverse real-world anomalies, shows promising results in detecting anomalies, demonstrating the effectiveness of CARLA in this domain.
+This research has demonstrated the potential of contrastive learning combined with synthetic anomaly injection to overcome the limitations of the lack of labelled data in TSAD.
 To enhance our model’s performance, future research will focus on refining positive and negative sample selection strategies.
-Moreover, we believe that there is much scope to improve performance by broadening the forms of injected anomalies used in our framework.
+Moreover, we believe that there is much scope to overcome the limitations of the anomaly injection component by broadening the forms of injected anomalies used in our framework.
 We further hypothesise that varying the severity of injected anomalies may improve the model’s ability to detect anomalies in representation space.
 We commend the investigation of these intriguing prospects in future research.
-Additionally, considering that some datasets with a significant number of abnormal samples, resulting in more anomalous windows during the Pretext Stage, it becomes critical to revisit the impact of selecting positives and negatives for abnormal anchor samples under unsupervised settings.
-This reconsideration is crucial for enhancing the model’s effectiveness in time series anomaly detection.
-In conclusion, this paper’s results, based on appropriate metrics, contribute significantly to further investigations and advancements in time series anomaly detection, particularly on datasets like WADI and KPI.
+Additionally, considering that some datasets have a significant number of abnormal samples, resulting in more anomalous windows during the Pretext Stage, it becomes critical to revisit the impact of selecting positives and negatives for abnormal anchors in unsupervised settings.
+This reconsideration is crucial for enhancing the model’s effectiveness in TSAD.
+Our results, based on appropriate metrics, contribute significantly to further investigations and advancements in TSAD, particularly on datasets like WADI and KPI.
 By employing the correct evaluation metrics, such as the conventional F1 score without point adjustment and AU-PR, this research provides a solid foundation for assessing the performance of future models in this field.
-The findings presented herein enable researchers to benchmark their models effectively and compare their results against established standards.
-Reliable evaluation metrics and validated results serve as valuable resources, driving progress and innovation in time series anomaly detection research.
+The findings presented herein enable researchers to benchmark their models and compare their results against established standards.
+Reliable evaluation metrics and validated results serve as valuable resources, driving innovation in TSAD research.
 Appendix A List of Symbols and Definitions
@@
 Symbol Definition
-$\mathcal{D}$
-Sequential time series windows where $\mathcal{D}=\{w_{1},w_{2},..,w_{m}\}$
-$w_{i}$
-$i^{th}$ window in $\mathcal{D}$ where $w_{i}=\{x_{1},x_{2},..,x_{WS}\}$
-$WS$
-Time series window size
-$Dim$
-Number of dimensions in time series $\mathcal{D}$
-$m$
-Number of windows
-$Q$
-Number of nearest/furthest neighbours
-$a$
-Anchor sample (i.e. the original time series window)
-$y$
-Number of temporally proximate windows
-$p$
-Positive sample (i.e. random window selected from $y$ temporally proximate previous windows)
-$n$
-Negative sample (i.e. anomaly injected window)
-$(a,p,n)$
-Triplet in Pretext Stage
-$\mathcal{T}$
-All triplets set in Pretext Stage
-$\alpha$
-Margin in pretext loss
-$\phi_{p}$
-Pretext trained model
-$\mathcal{B}$
-All neighbours set
-$\mathcal{N}$
-$Q$-Nearest neighbours set
-$\mathcal{F}$
-$Q$-Furthest neighbours set
-$\phi_{s}$
-Self-supervised trained model
-$C_{m}$
-Majority class label
-$\beta$
-Entropy loss weight
-Appendix B Point Adjustment (PA)
-A notable issue with many existing Time Series Anomaly Detection (TSAD) methods is their reliance on a specific evaluation approach called point adjustment (PA) strategy (Xu et al., 2018) and utilised in TSAD literature such as Su et al.
-(2019).
-PA operates under the following principle: if a given segment is detected as anomalous, all previous segments in that anomaly subsequence are considered as anomalies (see Figure 9).
-PA was introduced under the assumption that a solitary alert during an anomaly phase is adequate to recover the system.
-It has evolved into a fundamental stage in the assessment of TSAD, with certain studies exclusively documenting $F1_{PA}$ without the conventional F1 score Kim et al.
-(2022).
-A higher $F1_{PA}$ is often interpreted as indicating better detection capability.
-Figure 9: Impact of point adjustment (PA) on anomaly detection in time series.
-The figure represents a toy example using a sequence of 12 samples.
-Presents the anomaly scores for each sample within the sequence.
-The second row depicts the outcomes of anomaly detection using a threshold of 0.5 applied to all samples’ anomaly scores.
-The third row provides the actual labels (ground truth) for the samples within the sequence.
-The fourth and final row presents the adjusted anomaly detection results using point adjustment with the ground truth.
-Evaluation metrics are calculated for both unadjusted and adjusted anomaly detection.
-In Figure 9, for the unadjusted detection, we have True Positives (TP): 2, True Negatives (TN): 3, False Positives (FP): 1, False Negatives (FN): 6, yielding a Recall of 0.25, Precision of 0.67, and F1 score of 0.36.
-In contrast, the adjusted detection (in bold for emphasis) resulted in TP: 7, TN: 3, FP: 1, and FN: 1, giving a significantly improved Recall of 0.875, Precision of 0.875, and F1 score of 0.875.
-This illustrates the substantial difference in evaluation results due to the application of PA in the anomaly detection process.
-It is evident that the point adjustment greatly enhances the performance of anomaly detection.
-Table 9: Precision ($Prec_{PA}$), recall ($Rec_{PA}$), and $F1_{PA}$ results with point adjustment (PA) for various models on multivariate time series datasets.
+$\mathcal{D}$ Sequential time series windows where $\mathcal{D}=\{w_{1},w_{2},..,w_{m}\}$
+$w_{i}$ $i^{th}$ window in $\mathcal{D}$ where $w_{i}=\{x_{1},x_{2},..,x_{WS}\}$
+$WS$ Time series window size
+$Dim$ Number of dimensions in time series $\mathcal{D}$
+$m$ Number of windows
+$Q$ Number of nearest/furthest neighbours
+$a$ Anchor sample (i.e. the original time series window)
+$y$ Number of temporally proximate windows
+$p$ Positive sample (i.e. random window selected from $y$ temporally proximate previous windows)
+$n$ Negative sample (i.e. anomaly injected window)
+$(a,p,n)$ Triplet in Pretext Stage
+$\mathcal{T}$ All triplets set in Pretext Stage
+$\alpha$ Margin in pretext loss
+$\phi_{p}$ Pretext trained model
+$\mathcal{B}$ All neighbours set
+$\mathcal{N}$ $Q$-Nearest neighbours set
+$\mathcal{F}$ $Q$-Furthest neighbours set
+$\phi_{s}$ Self-supervised trained model
+$C_{m}$ Majority class label
+$\beta$ Entropy loss weight
+Table 9: Precision ($Prec_{PA}$), recall ($Rec_{PA}$), and $F1_{PA}$ results with point adjustment (PA) for various models on multivariate and univariate time series datasets.
+(a) Multivariate Time Series Datasets
 Model Metric MSL SMAP SMD SWaT WADI
@@
 $F1_{PA}$ 0.8919 0.7106 0.5353 0.8709 0.5138
-AnomalyTransformer $Prec_{PA}$ 0.7201 0.6481 0.8058 0.9087 0.0642
+AnomalyTran $Prec_{PA}$ 0.7201 0.6481 0.8058 0.9087 0.0642
 $Rec_{PA}$ 0.9832 0.9258 0.8482 0.7735 1.0000
@@
 $F1_{PA}$ 0.9064 0.8444 0.7515 0.8288 0.3122
-Table 10: Precision ($Prec_{PA}$), recall ($Rec_{PA}$), and $F1_{PA}$ results with point adjustment (PA) for various models on univariate time series datasets.
+(b) Univariate Time Series Datasets
 Model Metric Yahoo-A1 KPI
@@
 $F1_{PA}$ 0.8549 0.5507
-However, PA carries a notable peril of inflating the performance of the model.
-In most of the TSAD models, an anomaly score is produced to indicate the degree of abnormality, and the model detects a given point as an anomaly when the score surpasses a predefined threshold.
-Incorporating PA, the predictions based on the model that randomly generated anomaly scores uniformly outperform all state-of-the-art models, as depicted in Tables 9 and 10.
-It becomes challenging to assert that a model with a superior $F1_{PA}$ performs better than a model with a lower $F1_{PA}$ when random anomaly scores have the potential to result in a high $F1_{PA}$ comparable to that of a proficient detection model.
-Experimental results in Tables 9 and 10 demonstrate that the random anomaly score model can outperform all state-of-the-art models, highlighting the limitations of relying solely on $F1_{PA}$ for model evaluation.
+Appendix B Point Adjustment (PA)
+A notable issue with many existing TSAD methods is their reliance on the point adjustment (PA) strategy [31], as used in TSAD literature like [27].
+PA operates on the principle that if a segment is detected as anomalous, all previous segments in that anomaly subsequence are also considered anomalies (see Figure 9).
+Incorporating PA, the Random model outperform all state-of-the-art models, as depicted in Tables 9(a) and 9(b).
+Figure 9: Impact of point adjustment (PA) on anomaly detection in time series.
+The figure represents a toy example with a sequence of 12 samples, presenting the anomaly scores for each sample.
+The second row depicts the outcomes of anomaly detection using a threshold of 0.5.
+The third row provides the actual labels (ground truth).
+The fourth row presents the adjusted anomaly detection results using point adjustment with the ground truth.
+Evaluation metrics are calculated for both unadjusted and adjusted anomaly detection.